The New TEX FAQ Your 100 Questions Answered version 1.9c, date 1994/12/06

The UK TEX Users Group Committee after the original maintained by Bobby Bodenheimer

January 20, 1995

Note

This document is a version of the FAQ article that was published as the December 1994 edition of the UKTUG magazine Baskerville (which wasn’t formatted like this). The intention is that its content should be available to be fed back into the normal comp.text.tex FAQ posting some time soon, but that work isn’t yet complete. The article is also to be available as a www page.

Contents

A Introduction 2

B The Background 3 1 What is TEX? . . . . . . . . . . . . . . . . . 3 2 How should I pronounce “TEX”? . . . . . . . 3 3 What is ? . . . . . . . . . . . . 3 4 What is L^{A}TEX? . . . . . . . . . . . . . . . . 3 5 How should I pronounce “L^{A}TEX”? “L^{A}TEX 2ε”? 3 6 Should I use plain TEX or L^{A}TEX? . . . . . . 3 7 What are the AMS packages (AMS-TEX, etc.)? 4 8 What is L^{A}MS-TEX? . . . . . . . . . . . . . 4 9 What is Eplain? . . . . . . . . . . . . . . . 4 10 What is Lollipop? . . . . . . . . . . . . . . 4 11 What is Texinfo? . . . . . . . . . . . . . . . 4 12 If TEX is so good, how come it’s free? …. 4 13 What is the future of TEX? . . . . . . . . . . 5 14 What are TUG and TUGboat? . . . . . . . . 5 15 Are there nationally-based user groups, too? . 5 16 TUG Technical Working Groups . . . . . . . 5

C Documentation and Help 6 17 Books on TEX and its relations . . . . . . . . 6 18 Where to find this article . . . . . . . . . . . 7 19 Mailing lists about TEX and its friends …. 7 20 BibTEX Documentation . . . . . . . . . . . 7 21 The PICTEX manual . . . . . . . . . . . . . 7 22 Finding TEX/L^{A}TEX macro packages . . . . . 7 23 Finding L^{A}TEX 2ε-related packages . . . . . . 7

D Bits and pieces of TEX 8 24 Virtual fonts . . . . . . . . . . . . . . . . . 8 25 \special commands . . . . . . . . . . . . 8 26 Documented L^{A}TEX sources (.dtx files) … 8 E Acquiring the Software 8 27 Repositories of TEX material . . . . . . . . . 8 28 Contributing a file to the archives . . . . . . 9 29 Finding new fonts . . . . . . . . . . . . . . 9 30 TEX CD-ROMs . . . . . . . . . . . . . . . . 9

F TEX Systems 9 31 TEX/L^{A}TEX for different machines . . . . . . . 9 32 TEX-friendly editors and shells . . . . . . . . 10 33 Commercial TEX implementations . . . . . . 10

G DVI Drivers and Previewers 12 34 DVI to PostScript conversion programs … 12 35 DVI drivers for HP LaserJet . . . . . . . . . 12 36 DVI previewers . . . . . . . . . . . . . . . 12

H Support Packages for TEX 12 37 Fig, a TEX-friendly drawing package . . . . . 12 38 TEXCAD, a drawing package for L^{A}TEX … 12 39 Spelling checkers for work with TEX . . . . . 12 40 The VORTEX package . . . . . . . . . . . . . 13

I Literate programming 13 41 What is Literate Programming? . . . . . . . 13 42 WEB for C, FORTRAN, and other languages 13

J Format conversions 13 43 Conversion between TEX/L^{A}TEX and others . . 13 44 Conversion from TEX/L^{A}TEX to plain ASCII . 14 45 Conversion from HTML or SGML to TEX . . 14 46 Conversion to HTML . . . . . . . . . . . . 14

K �������� 14 47 Getting to do what you want . . 14 48 Which font files should be kept . . . . . . . . 15 49 Getting bitmaps from the archives . . . . . . 15 L PostScript and TEX 15

50 Using PostScript fonts with TEX . . . . . . . 15 51 Previewing files using PostScript fonts …. 16 52 TEX font metric files for PostScript fonts … 16 53 Problems using PostScript fonts . . . . . . . 16 54 Choice of scalable outline fonts . . . . . . . 17 55 Including a PostScript figure in L^{A}TEX …. 18

M Special sorts of typesetting 18 56 Double-spaced documents in L^{A}TEX . . . . . 18 57 Formatting a thesis in L^{A}TEX . . . . . . . . . 18 58 Including a file in verbatim in L^{A}TEX . . . . . 18 59 Generating an index in TEX/L^{A}TEX . . . . . . 18 60 Using BibTEX with plain TEX . . . . . . . 19 61 Typesetting music in TEX . . . . . . . . . . 19 62 Drawing Feynman diagrams in L^{A}TEX …. 19

N How do I do X in TEX or L^{A}TEX 19 63 Proof environment . . . . . . . . . . . . . . 19 64 Symbols for the number sets . . . . . . . . . 19 65 Roman theorems . . . . . . . . . . . . . . . 19 66 Labels on lists . . . . . . . . . . . . . . . . 19 67 Unnumbered sections in the Table of Contents 20 68 Style of section headings . . . . . . . . . . . 20 69 Indent after section headings . . . . . . . . . 20 70 Changing the margins in L^{A}TEX . . . . . . . . 20 71 Insisting on changing the margins in L^{A}TEX . . 20 72 Finding the width of a letter, word, or phrase . 21 73 Excluding blocks of text from the .dvi file . 21 74 Defining a new log-like function in L^{A}TEX . . 21 75 Typesetting all those TEX-related logos …. 21 O Things are Going Wrong... 21

76 Weird hyphenation of words . . . . . . . . . 21 77 (Merely) peculiar hyphenation . . . . . . . . 21 78 Enlarging TEX . . . . . . . . . . . . . . . . 22 79 Moving tables and figures in L^{A}TEX . . . . . . 22 80 \pagestyle{empty} on first page in L^{A}TEX . 22 81 Odd behaviour of \rm, \bf, etc. . . . . . . . 22 82 Old L^{A}TEX font references such as \tenrm . . 22 83 Missing symbols . . . . . . . . . . . . . . . 23 84 L^{A}TEX gets cross-references wrong . . . . . . 23 85 \@ and @ in macro names . . . . . . . . . . . 23 86 Where are the msx and msy fonts? . . . . . . 23 87 Where are the am fonts? . . . . . . . . . . . 23 88 ‘String too long’ in BibTEX . . . . . . . . . 23

P Why does it do that? 24 89 Why does it ignore paragraph parameters? . . 24 90 What’s the reason for ‘protection’? . . . . . . 24 91 Why doesn’t \verb work within...? . . . . . 24 92 Case-changing oddities . . . . . . . . . . . . 24 93 Why are # signs doubled in macros? . . . . . 24

Q Recent Developments 25 94 The New Font Selection Scheme (NFSS) . . 25 95 L^{A}TEX 2ε (the new standard L^{A}TEX) . . . . . . 25 96 The L^{A}TEX3 project . . . . . . . . . . . . . . 25 97 Using TEX to typeset from SGML files …. 25 98 The Omega project . . . . . . . . . . . . . . 25 99 The NTS project . . . . . . . . . . . . . . . 25

R Perhaps There isn’t an Answer 26 100 What to do if you find a bug . . . . . . . . . 26

A Introduction

This article was prepared by the Committee of the UK TEX Users Group (UK TUG)^{1} as a development of a regular posting to the Usenet newsgroup comp.text.tex that has been maintained for some time by Bobby Bodenheimer (bobby@hot.

caltech.edu).

Usenet is a mechanism for exchanging articles between people who share interests or needs^{2}; a newsgroup is an area within Usenet carrying a particular class of articles. Since a common sort of article asks for help, advice or information, and since certain of these questions are regularly repeated (often with monotonous regularity), some public-spirited souls took to writing articles which listed “Frequently Asked Questions” and answers to them. Many members of UK TUG do not have access to Usenet, but could be expected to value the answers about TEX that have accumulated over the years; so we decided to update the list and publish it in Baskerville; we are grateful to Bobby for his permission to use his article in this way. As a quid pro quo, we are making the source of the article freely available, and it can be compiled by anyone who runs a production L^{A}TEX 2ε (see question 95), and has the required fonts. It is the committee’s hope that it will also be possible for the content of this article to feed back to the world-wide TEX community via Bobby’s regular posting.

In addition, a translation of the article is available on the

World-Wide Web, via URL http://www.cogs.susx.ac.
uk/cgi-bin/texfaq2html?introduction=yes^{3}

We have rearranged Bobby’s article quite a lot, and have added new questions and answers on the basis of our experience of answering questions about TEX, writing documents in TEX, and developing macros for TEX, over the years. We have also pruned it to take account of the changes that have happened in the world of TEX since Bobby first started.

The committee is grateful for help and advice, from the following outside its number: Barbara Beeton, Karl Berry, Damian Cugley, Michael Downes, Berthold Horn, Werner Icking, Ted Nieland, Pat Rau, Joachim Schrod, Philip Taylor, Ulrik Vieth, Rick Zaccone and Reinhard Zierke.

Finding the Files

Unless otherwise specified, all files mentioned in this article are available from a CTAN archive, or from one of their mirrors.

1For 1993-94: Peter Abbott, Rosemary Bailey, Malcolm Clark, Robin Fairbairns, Jonathan Fine, Carol Hewlett, Alan Jeffrey, Sebastian Rahtz, and Chris Rowley; in addition, for 1994–95, David Carlisle

2Usenet, as its name implies, is a means of using some sort of network; in the earliest days the network was made by stringing together a series of telephone lines, but nowadays Usenet is most often carried over the Internet

3This is a temporary URL; a final home for the document is to be provided in due course

Question 27 gives details of the CTAN archives, and how to retrieve files from them. If you don’t have access to the internet, question 30 tells you of sources of CD-ROMs that offer snapshots of the archives.

The reader should also note that the first directory name of the path name of every file on CTAN has been elided from what follows, for the simple reason that it’s always the same

(tex-archive/).

To avoid confusion, we’ve elided the full stop^{4} from the end of any sentence whose last item is a path name (note that such sentences only occur at the end of paragraphs). Though the path names are set in a different font from running text, it’s not easy to distinguish the font of a single dot!

B The Background

1 What is TEX?

TEX is a typesetting system written by Donald E. Knuth, who says in the Preface to his book on TEX (see question 17) that it is “intended for the creation of beautiful books—and especially for books that contain a lot of mathematics”.

Knuth developed a system of ‘literate programming’ to write TEX, and he provides the literate (WEB) source of TEX free of charge, together with tools for processing the web source into something that can be compiled and something that can be printed; there’s never any mystery about what TEX does. Furthermore, the WEB system provides mechanisms to port TEX to new operating systems and computers; in order that one may have some confidence in the ports, Knuth supplied a test by means of which one may judge the fidelity of a TEX system. TEX and its documents are therefore highly portable.

TEX is a macro processor, and offers its users a powerful programming capability. For this reason, TEX on its own is a pretty difficult beast to deal with, so Knuth provided a package of macros for use with TEX called plain TEX; plain TEX is effectively the minimum set of macros one can usefully employ with TEX, together with some demonstration versions of higher-level commands (the latter are better regarded as models than used as-is). When people say they’re “programming in TEX”, they usually mean they’re programming in plain TEX.

2 How should I pronounce “TEX”?

The ‘X’ stands for the Greek letter Chi (χ), and is pronounced by English-speakers either a bit like the ‘ch’ in ‘loch’ ([x] in the IPA) or like ‘k’. It definitely is not pronounced ‘ks’.

3 What is ��������?

was written by Knuth as a companion to TEX; whereas TEX defines the layout of glyphs on a page, - defines the shapes of the glyphs and the relations between them. details the sizes of glyphs, for TEX’s benefit, and details the rasters used to represent the glyphs, for the benefit of programs that will produce printed output as post processes after a run of TEX.

’s language for defining fonts permits the expression of several classes of things: first (of course), the simple geometry of the glyphs; second, the properties of the print engine for which the output is intended; and third, ‘meta’- information which can distinguish different design sizes of the same font, or the difference between two fonts that belong to the same (or related) families.

Knuth (and others) have designed a fair range of fonts using , but font design using is much more of a minority skill than is TEX macro-writing. The complete TEX-user nevertheless needs to be aware of , and to be able to run to generate personal copies of new fonts.

4 What is L^{A}TEX?

L^{A}TEX is a TEX macro package, written by Leslie Lamport, that provides a document processing system. L^{A}TEX allows markup to describe the structure of a document, so that the user need not think about presentation. By using document classes and addon packages, the same document can be produced in a variety of different layouts.

Lamport says that L^{A}TEX “represents a balance between functionality and ease of use”. This shows itself as a continual conflict that leads to the need for such as the present article: L^{A}TEX can meet most user requirements, but finding out how is often tricky.

5 How should I pronounce “L^{A}TEX”? “L^{A}TEX 2ε”?

Lamport never recommended how one should pronounce L^{A}TEX, but a lot of people pronounce it ‘Lay TEX’ or perhaps ‘Lah TEX’ (with TEX pronounced as the program itself; see question 2).

The ‘epsilon’ in ‘L^{A}TEX 2ε’ is supposed to be suggestive of a small improvement over the old L^{A}TEX 2.09. Nevertheless, most people pronounce the name as ‘L^{A}TEX-two-ee’.

6 Should I use plain TEX or L^{A}TEX?

There’s no straightforward answer to this question. Many people swear by plain TEX, and produce highly respectable documents using it (Knuth is an example of this, of course). But equally, many people are happy to let someone else take the design decisions for them, accepting a small loss of flexibility in exchange for a saving of brain power.

The arguments around this topic can provoke huge amounts of noise and heat, without offering much by way of light; your best bet is to find out what those around you are using, and to go with the crowd. Later on, you can always switch your allegiance; don’t bother about it.

If you are preparing a manuscript for a publisher or journal, ask them what markup they want before you develop your own; many big publishers have developed their own L^{A}TEX styles for journals and books, and insist that authors stick closely to their markup.

4‘Full stop’ (British English)==‘period’ (American English)

7 What are the AMS packages (AMS-TEX, etc.)?

AMS-TEX is a TEX macro package, originally written by Michael Spivak for the American Mathematical Society (AMS) during 1983–1985. It is described in The Joy of TEX by Michael D. Spivak (second edition, AMS, 1990, ISBN 0-821-82997-1). It is based on plain TEX, but provides many features for producing more professional-looking maths formulas with less burden on authors. It pays attention to the finer details of sizing and positioning that mathematical publishers care about. The aspects covered include multi-line displayed equations, equation numbering, ellipsis dots, matrices, double accents, multiline subscripts, syntax checking (faster processing on initial error-checking TEX runs), and other things.

As L^{A}TEX increased in popularity, authors asked to submit papers to the AMS in L^{A}TEX, and so the AMS developed AMS-L^{A}TEX, which is a collection of L^{A}TEX packages and classes that offer authors most of the functionality of AMS-TEX.

8 What is L^{A}_{M}S-TEX?

L^{A}MS-TEX was Michael Spivak’s fusion of AMS-TEX and L^{A}TEX. Its strong points are:

• a commutative diagram package that produces very pleasing output;

• a separate program dvipaste for producing complex tables separately from the main document (thus avoiding problems from exceeding TEX’s main memory capacity); and

• extensive control at the user level over the formatting of automatically-generated numbers.

However, AMS-L^{A}TEX had come out by the time L^{A}MS-TEX was released, so that L^{A}MS-TEX never really caught on.

9 What is Eplain?

The Eplain macro package expands on and extends the definitions in plain TEX. Eplain is not intended to provide “generic typesetting capabilities”, as do L^{A}TEX or Texinfo (see question 11). Instead, it provides definitions that are intended to be useful regardless of the high-level commands that you use when you actually prepare your manuscript.

For example, Eplain does not have a command \section, which would format section headings in an “appropriate” way, as L^{A}TEX’s \section. The philosophy of Eplain is that some people will always need or want to go beyond the macro designer’s idea of “appropriate”. Such canned macros are fine — as long as you are willing to accept the resulting output. If you don’t like the results, or if you are trying to match a different format, you are out of luck.

On the other hand, almost everyone would like capabilities such as cross-referencing by labels, so that you don’t have to put actual page numbers in the manuscript. Karl Berry, the author of Eplain, says he is not aware of any generally available macro packages that do not force their typographic style on an author, and yet provide such capabilities.

10 What is Lollipop?

Lollipop is a macro package written by Victor Eijkhout; it was used in the production of his book “TEX by Topic” (see question 17). The manual says of it:

Lollipop is ‘TEX made easy’. Lollipop is a macro package that functions as a toolbox for writing TEX macros. It was my intention to make macro writing so easy that implementing a fully new layout in TEX would become a matter of less than an hour for an average document, and that it would be a task that could be accomplished by someone with only a very basic training in TEX programming.

Lollipop is an attempt to make structured text formatting available for environments where previously only wysiwyg packages could be used because adapting the layout is so much more easy with them than with traditional TEX macro packages.

The manual goes on to talk of ambitions to “capture some of the L^{A}TEX market share”; it’s a very witty package, but little sign of it taking over from L^{A}TEX is detectable... An article about Lollipop appeared in TUGboat 13(3), 1992.

11 What is Texinfo?

Texinfo is a documentation system that uses one source file to produce both on-line information and printed output. So instead of writing two different documents, one for the on-line help and the other for a typeset manual, you need write only one document source file. When the work is revised, you need only revise one document. You can read the on-line information, known as an “Info file”, with an Info documentation-reading program. By convention, Texinfo source file names end with a .texi or .texinfo extension. You can write and format Texinfo files into Info files within GNU emacs, and read them using the emacs Info reader. If you do not have emacs, you can format Texinfo files into Info files using makeinfo and read them using info.

A set of TEX macros for formatting Texinfo files are avail-

                                                     able in macros/texinfo/texinfo.tex

12 If TEX is so good, how come it’s free?

It’s free because Knuth chose to make it so. He is nevertheless apparently happy that others should earn money by selling TEX-based services and products. While several valuable TEXrelated tools and packages are offered subject to restrictions imposed by the GNU General Public Licence (‘Copyleft’), TEX itself is not subject to Copyleft.

There are commercial versions of TEX available; for some users, it’s reassuring to have paid support. What is more, some of the commercial implementations have features that are not available in free versions. (The reverse is also true: some free implementations have features not available commercially).

Usually, this article does not describe commercial versions. Question 33 lists the major vendors.

13 What is the future of TEX?

Knuth has declared that he will do no further development of TEX; he will continue to fix any bugs that are reported to him (though bugs are rare). This decision was made soon after TEX version 3.0 was released; at each bug-fix release the version number acquires one more digit, so that it tends to the limit π. Knuth wants TEX to be frozen at version π when he dies; thereafter, no further changes may be made to Knuth’s source. (A similar rule is applied to ; its version number tends to the limit e.)

There are projects (some of them long-term projects: see, for example, questions 96 and 97) to build substantial new macro packages based on TEX. For the even longer term, there are various projects to build a successor to TEX; see questions 98 and 99.

14 What are TUG and TUGboat?

TUG is the TEX Users Group. TUGboat is TUG’s main journal, containing useful articles about TEX and . TUG also produces a newsletter for members (TEX and TUG News), organises a yearly conference, runs training courses, sells almost all TEX-related books, and distributes TEX-related microcomputer software on disk. TUG has a Technical Council to coordinate TEX-related developments (see see question 16). Enquiries should be directed to:

TEX Users Group P. O. Box 869 Santa Barbara, CA 93102-0869 USA Tel: (+1) 805-963-1338 Fax: (+1) 805-963-8358

Email: tug@tug.org
CTAN details: usergrps/tug

15 Are there nationally-based user groups, too?

The following groups publish their membership (etc.) information electronically on CTAN archives:

DANTE, Deutschsprachige Anwendervereinigung TEX e.V. Postfach 10 18 40 D-69008 Heidelberg Germany Tel: (+49) 06221/2 97 66 Fax: (+49) 06221/16 79 06

Email: dante@dante.de
CTAN details: usergrps/dante

GUTenberg, BP 10, 93220 Gagny principal, France

CTAN details: usergrps/gut

NTG Postbus 394, 1740AJ Schagen, The Netherlands

Email: ntg@nic.surfnet.nl
CTAN details: usergrps/ntg

UK TEX Users’ Group, c/o Peter Abbott, 1 Eymore Close, Selly Oak, Birmingham B29 4LB UK Tel: (+44) 0121 476 2159

Email: UKTuG-Enquiries@tex.ac.uk
CTAN details: usergrps/uktug

A listing of all known groups is available as usergrps/

info/usergrps.tex

16 TUG Technical Working Groups

TUG (see question 14) has an autonomous Technical Council which oversees a number of working groups on areas of common interest to the TEX community. The Council has three members (current chair is Michael Ferguson, assisted by Yannis Haralambous and Sebastian Rahtz), who liaise with chair people of each working group. Each group establishes its own working methods and membership, and anyone interested in taking part should contact the chair. Suggestions for new groups should be addressed to Michael Ferguson (mike@

inrs-telecom.uquebec.ca).

A brief list of the active groups follows:

WG-92-00 (IRP-TWG) Independent Research Project TWG. To recognise and report to the TEX Board and the TEX Community on important projects which are independent of TUG but are of concern to the entire TEX Community.

Contact: Alan Hoenig (ajhjj@cunyvm.cuny.edu)

WG-92-01 TEX Extended Mathematic Font Encoding. To create font encoding standards for Mathematical fonts used in TEX systems.

Contact: Barbara Beeton (bnb@math.ams.org)

WG-92-03 Multiple Language Coordination.

The primary purpose of this working group is to obtain, for TEX systems, a consistent means for implementing, accessing, and describing, the fonts, ligature rules, hyphenation patterns and other special requirements for a given linguistic group.

Contact: Yannis Haralambous (Yannis.Haralambous@

univ-lille1.fr)

WG-92-04 TEX for the Disabled.

The primary purpose of this working group is as a forum for those people interested in using and/or enhancing TEX to serve the needs of those with visual and other disabilities.

Contact: T.V. Raman (raman@crl.dec.com)

WG-92-05 TEX Archive Guidelines.

The purpose of this Technical Working Group is to develop guidelines for the effective management and utilisation of major TEX archives, and to initiate communication among the maintainers of the existing archives for the purpose of coordination and synchronisation.

Contact: George Greenwade (bed_gdg@shsu.edu)

WG-94-07 TEX Directory Structures.

The primary purpose of this TWG is to identify a universal directory structure for macros, fonts and other related TEX software so that recommendations can be made to all suppliers of TEX software.

Contact: Norm Walsh (norm@ora.com)

WG-94-08 DVI Driver Implementation and Standardisation Issues. The major objective shall be to study the issues in the requirements of DVI Drivers imposed by changing needs and technologies, and to make recommendations for implementation and standardisation of such drivers to enhance the uniformity of their use. Work will include, but not be limited to, the examination of the use, syntax, and semantics of \special{..} commands.

Contact: Michael Sofka (sofkam@rpi.edu)

WG-94-09 TEX and SGML.

The major objective is to investigate the requirements and difficulties in developing an interface technology for TEX and SGML.

Contact: Ken Dreyhaupt (kend@springer-ny.com)

WG-94-10 TEX and Linguistics.

The main goal is to study and discuss the requirements for typesetting linguistics in TEX and as a means of identifying, examining, testing, and comparing macros, fonts, style files and other aids for typesetting linguistics.

Contact: Christina Thiele (cthiele@ccs.carleton.

ca)

C Documentation and Help

17 Books on TEX and its relations

While Knuth’s book is the definitive reference for TEX, there are other books covering TEX:

The TEXbook by Donald Knuth (Addison-Wesley, 1984, ISBN 0-201-13447-0, paperback ISBN 0-201-13448-9)

A Beginner’s Book of TEXby Raymond Seroul and Silvio Levy, (Springer Verlag, 1992, ISBN 0-387-97562-4)

Introduction to TEX by Norbert Schwarz (Addison-Wesley, 1989, ISBN 0-201-51141-X)

A Plain TEX Primer by Malcolm Clark (Oxford University Press, 1993, ISBNs 0-198-53724-7 (hardback) and 0-198-53784-0 (paperback))

TEX by Topic by Victor Eijkhout (Addison-Wesley, 1992, ISBN 0-201-56882-9)

TEX for the Beginner by Wynter Snow (Addison-Wesley, 1992, ISBN 0-201-54799-6)

TEX for the Impatient by Paul W. Abrahams, Karl Berry and Kathryn A. Hargreaves (Addison-Wesley, 1990, ISBN0-201-51375-7)

TEX in Practice by Stephan von Bechtolsheim (Springer Verlag, 1993, 4 volumes, ISBN 3-540-97296-X for the set, or Vol. 1: 0-387-97595-0, Vol. 2: 0-387-97596-9, Vol. 3: 0-387-97597-7, and Vol. 4: 0-387-97598-5)

TEX: Starting from 1 ^{5} by Michael Doob (Springer Verlag, 1993, ISBN 3-540-56441-1)

For L^{A}TEX, see:

L^{A}TEX, a Document Preparation System by Leslie Lamport (second edition, Addison Wesley, 1994, ISBN 0-201-15790-X)

A guide to L^{A}TEXHelmut Kopka and Patrick W. Daly (Addison-Wesley, ISBN 0-201-56889-6)

The L^{A}TEX Companion by Michel Goossens, Frank Mittelbach, and Alexander Samarin (Addison-Wesley, 1993, ISBN 0-201-54199-8)

L^{A}TEX Notes: Practical Tips for Preparing Technical Documents by J. Kenneth Shultis (Prentice Hall, 1994, ISBN 0-131-20973-6)

L^{A}TEX Line by Lineby Antoni Diller (John Wiley & Sons, 1993, ISBN 0-471-93471-2)

L^{A}TEX for Scientists and Engineers by David J. Buerger (McGraw-Hill, 1990, ISBN 0-070-08845-4)

Math into TEX: A Simplified Introduction using AMS-L^{A}TEX by George Grätzer (Birkhäuser, 1993, ISBN 0-817-63637-4, or, in Germany, ISBN 3-764-33637-4)

Of that list, only Lamport, and Goossens, Mittelbach and Samarin cover L^{A}TEX 2ε, though at least some new editions are reported to be in preparation.

The list for is rather short:

The book by Donald Knuth (Addison Wesley, 1986, ISBN 0-201-13445-4)

A book covering a wide range of topics (including installation and maintenance) is:

Making TEX Workby Norman Walsh (O’Reilly and Associates, Inc, 1994, ISBN 1-56592-051-1)

5That’s ‘Starting from Square One’

18 Where to find this article

The article, on which the present one is based, is posted monthly to newsgroup comp.text.tex and cross-posted to

newsgroups news.answers and comp.answers. The most

recently posted copy of the article is at CTAN in info/ TeX-FAQ; it is also archived at any site that archives news. answers, such as rtfm.mit.edu (18.181.0.24), and this article is available there via anonymous ftp (in the directory

pub/usenet/news.answers/tex-faq). If you have access

to email, but not to ftp, send an e-mail message containing the

lines “SENDME FAQ” to fileserv@shsu.edu.

A version of the article may be browsed via the World-Wide

Web, at URL http://www.cogs.susx.ac.uk/cgi-bin/
texfaq2html?introduction=yes^{6}

19 Mailing lists about TEX and its friends

There are (still) people who can use networks but can’t read Usenet news; for them, not all is lost if they can send and receive email.

The UKTEX digest and the TEXhax digest are both operated as mailing lists. Send a message ‘subscribe uktex’

to uktex@tex.ac.uk, or ‘subscribe texhax’ to texhax@

tex.ac.uk to join them.

The (rather high volume) Usenet group comp.text.tex is gatewayed to the info-tex mailing list; subscribe by

sending a message ‘subscribe info-tex <your name>’ to

listserv@shsu.edu. The group is also available in digest form through the ctt-digest list; subscribe by send-

ing a message ‘subscribe ctt-Digest <your name>’ to
listserv@shsu.edu; ctt-Digest is approximately the same

size as info-tex, but it comes in fewer messages, which is an advantage on some systems.

Several TEX-related lists may be accessed via listserv@

vm.urz.uni-heidelberg.de. Send a message containing

the line “help” to this address.

The literate programming newsgroup (see question 41)

comp.programming.literate is gatewayed to the

litprog mailing list; subscribe by sending a message

‘subscribe litprog <your name>’ to listserv@shsu.
edu.

20 BibTEX Documentation

BibTEX, a program originally designed to produce bibliographies in conjunction with L^{A}TEX, is explained in Section 4.3 and Appendix B of Leslie Lamport’s L^{A}TEX manual (see question 17). The document “BibTEXing”, contained in the file btxdoc.tex, gives a more complete description. The L^{A}TEX Companion (see question 17) also has information on BibTEX and writing BibTEX style files.

The document “Designing BibTEX Styles”, contained in the file btxhak.tex, explains the postfix stack-based language used to write BibTEX styles (.bst files). The file btxbst.doc is the template for the four standard styles (plain, abbrv, alpha, unsrt). It also contains their documentation. The complete BibTEX documentation set (including the files above) is

in biblio/bibtex/distribs/doc

There is a Unix BibTEX man page in the web2c package (see question 31). Throw away any copy you may find of an obsolete man page written in 1985 before “BibTEXing” and “Designing BibTEX Styles” appeared.

21 The PICTEX manual

PICTEX is a set of macros for drawing diagrams and pictures. The macros are freely available in graphics/pictex; however, the PICTEX manual itself is not free. It is available for $30 ($35 with the disk) from the TEX Users Group (see question 14). The proceeds from the sales go to Michael Wichura, the author of PICTEX, and to TUG.

22 Finding TEX/L^{A}TEX macro packages

Before you ask for a TEX macro or L^{A}TEX class or package file to do something, try searching the TEX macro index written by

                                                     David M. Jones (dmjones@theory.lcs.mit.edu) and avail-
                                                     able in help/Tex-index/TeX-index

The index is an excellent reference document with plenty of cross-references, but it is somewhat out of date. Until the present work to bring the index up to date is complete, you are best advised to ignore any file location information in it; having learnt of a file that seems interesting, search a CTAN archive for it (see below for the simplest method).

To use the searching features of the CTAN archives (see question 27), establish an anonymous ftp connection to a CTAN site, and type the command ‘quote site index <term>’. The archive will provide a list of files with the string <term> in their names^{7}. (Some ftp clients don’t implement the command ‘quote’, in most such cases, you need simply to type ‘site index <term>’.)

23 Finding L^{A}TEX 2ε-related packages

A common confusion that can arise is that no .sty or .cls file exists on the CTAN archives for the package or class you’ve been recommended to acquire. For example, there are many packages listed in The L^{A}TEX Companion (see question 17) without details of their location on CTAN. The file info/companion.ctan may be of help in finding such files, but it’s not a general solution.

The canonical way of distributing L^{A}TEX 2ε packages and classes is as .dtx files (see question 26). Of course, if your site index command was for ‘package.sty’, you won’t find anything — you should simply ask for ‘package’ (the search can find things on the basis of substrings of their name). In the case that what you’re seeking is indeed distributed this way, you should retrieve the whole directory that the .dtx file is in, which will get you all the related files, including one or more .ins (installation) files. Process the ‘appropriate’ .ins file using L^{A}TEX 2ε, and you will generate the file (or files) that you want.

6This is a temporary URL; a final home for the document is to be provided in due course

7In fact, <term> is an Unix regular expression; the command is searching an index file using the command grep

D Bits and pieces of TEX

24 Virtual fonts

Virtual fonts for TEX were first implemented by David Fuchs in the early days of TEX, but for most people they started when Knuth redefined the format, and wrote some support software, in 1989. Virtual fonts provide a way of telling TEX about something more complicated than just a one-to-one character mapping. The entities you define in a virtual font look like characters to TEX (they appear with their sizes in a font metric file), but the dvi processor may expand them to something quite different. You can use this facility just to remap characters, to make a composite font with glyphs drawn from several sources, or to build up an effect in arbitrarily complicated ways — a virtual font may contain anything which is legal in a dvi file. In practice, the most common use of virtual fonts is to remap PostScript fonts (see question 52) or to build ‘fake’ maths fonts.

It is important to realise that TEX itself does not see virtual fonts; for every virtual font read by the dvi driver there is a corresponding tfm file read by TEX. Virtual fonts are normally created in a single ASCII vpl (Virtual Property List) file, which includes both sets of information. The vptovf program is then used to the create the binary tfm and vf files. The commonest way (nowadays) of generating vpl files is to use the fontinst package, which is described in detail in question 52. fonts/ utilities/qdtexvpl is another utility for creating ad-hoc virtual fonts.

25 \special commands

TEX provides the means to express things that device drivers can do, but about which TEX itself knows nothing. For example, TEX itself knows nothing about how to include PostScript figures into documents, or how to set the colour of printed text; but some device drivers do.

Such things are introduced to your document by means of \special commands; all that TEX does with these commands is to expand their arguments and then pass the command to the dvi file. In most cases, there are macro packages provided (often with the driver) that provide a comprehensible interface to the \special; for example, there’s little point including a figure if you leave no gap for it in your text, and changing colour proves to be a particularly fraught operation that requires real wizardry. L^{A}TEX 2ε has standard graphics and colour packages that make file inclusion, rotation, scaling and colour via \specials all easy.

The allowable arguments of \special depend on the device driver you’re using. Apart from the examples above, there are \special commands in the emTEX drivers (e.g., dvihplj, dviscr, etc.) that will draw lines at arbitrary orientations, and commands in dvitoln03 that permit the page to be set in landscape orientation.

26 Documented L^{A}TEX sources (.dtx files)

L^{A}TEX 2ε, and many support macro packages, are now written in a literate programming style (see question 41), with source and documentation in the same file. This format, known as ‘doc’, was originated by Frank Mittelbach. The documented sources conventionally have the suffix .dtx, and should normally be stripped of documentation before use with L^{A}TEX. Alternatively you can run L^{A}TEX on a .dtx file to produce a nicely formatted version of the documented code. An installation script (with suffix .ins) is usually provided, which needs the standard L^{A}TEX 2ε docstrip package (among other things, the installation process strips all the comments that make up the documentation for speed when loading the file into a running L^{A}TEX system). Several packages can be included in one .dtx file, with conditional sections, and there facilities for indices of macros etc. Anyone can write .dtx files; the format is explained in The L^{A}TEX Companion (see question 17). There are no programs yet to assist in composition.

.dtx files are not used by L^{A}TEX after they have been processed to produce .sty or .cls (or whatever) files. They need not be kept with the working system; however, for many packages the .dtx file is the primary source of documentation, so you may want to keep .dtx files elsewhere.

E Acquiring the Software

27 Repositories of TEX material

To aid the archiving and retrieval of of TEX-related files, a TUG working group developed the Comprehensive TEX Archive Network (CTAN). Each CTAN site has identical material, and maintains authoritative versions of its material. These collections are extensive; in particular, almost everything mentioned in this article is archived at the CTAN sites, even if its location isn’t explicitly stated.

The CTAN sites are currently ftp.dante.de (129.206.100.192), ftp.tex.ac.uk (128.232.1.87), and ftp. shsu.edu (192.92.115.10). The organisation of TEX files on all these sites is identical and starts at tex-archive/. To reduce network load, please use the CTAN site or mirror closest to you. A complete and current list of CTAN sites and mirrors can be obtained by using the finger utility on ‘user’ ctan_us@ ftp.SHSU.edu; it is also available as file CTAN.sites

To find software at a CTAN site, use anonymous ftp to the host, and then execute the command ‘quote site index <term>’ (see question 22 for details).

                                                        Email servers ftpmail@ftp.shsu.edu and ftpmail@

ftp.dante.de provide an ftp-like interface through mail. Send a message containing just the line ‘help’ to your nearest server for details of use.

Users on BITNET may access anonymous ftp for some files can be had indirectly by sending mail to BITFTP@PUCC. BITNET. Send a message containing the line “help” to this address for more information.

There is also the DECUS TEX collection of material for VMS, Unix, MS-DOS, and the Macintosh. The material for VMS has not been kept up-to-date, but continues to run on OpenVMS on the VAX. It is available via anonymous ftp from

                                                     wuarchive.wustl.edu (128.252.135.4) in decus/tex/. It

can also be obtained from the DECUS Library (reference number VS0058) in the US, or through your DECUS office outside of the US. To contact the DECUS Library, send mail or telephone:

DECUS

LIBRARY ORDER PROCESSING 334 South Street, SHR3-1/T25 Shrewsbury, MA 01545-4195 USA Tel: 800-DECUS55 (within the USA, for information) Fax: (+1) 508-841-3373 (for inquiries)

or send electronic mail for information to the DECUS TEX Collection Editor, Ted Nieland (nieland@ted.hcst.com).

Finally, of course, the TEX user who has no access to any sort of network may buy a copy of the archive on CD-ROM (see question 30).

28 Contributing a file to the archives

Use anonymous ftp to any CTAN archive (see question 27) and retrieve the file README.uploads in the root directory. It contains instructions for uploading files and notifying the appropriate people for that site.

If you cannot use ftp, mail your contribution to sty-mgr@ shsu.edu and it will be passed along. You will make everyone’s life easier if you choose a descriptive and unique name for your submission, so it’s probably a good idea to check that your style file’s name is not already in use by means of the ‘site index’ command (see question 27).

29 Finding new fonts

A comprehensive list of fonts is posted to comp. fonts and to comp.text.tex, roughly every six weeks, by Lee Quin (lee@sq.sq.com); it is available as info/

metafont-list

The list contains details both of commercial fonts and of fonts available via anonymous ftp. Most of the fonts are available via anonymous ftp from the CTAN archives (see question 27).

30 TEX CD-ROMs

If you don’t have access to the Internet, you can get the CTAN collections on a CD-ROM. Even those who do will find it very convenient to have 600Mb of TEX-related files to hand. Prime Time Freeware produced TEXcetera 1.1 in July 1994, which is a snapshot of CTAN taken in June 1994. Regular updates are planned. The material is all compressed in ZIP format to fit it all on one CD, and to avoid the limitations of the ISO 9660 file system directory. You can buy the CD from:

Prime Time Freeware 370 Altair Way, Suite 150 Sunnyvale CA 94086 USA Tel: (+1) 408 433 9662 Fax: (+1) 408 433 0727

Email: ptf@cfcl.com

or from many CD-ROM resellers, or the TUG office (see question 14). Price will be around $60. Please note that PTF is not a big commercial firm, and is a good friend of the TEX community.

If you want a ready-to-run TEX system on CD-ROM, one is available for DOS only (so far). The Dutch TEX Users Group (NTG) publish the whole 4AllTEX workbench on a CD-ROM packed with all the MS-DOS TEX software, macros and fonts you can want. It is available from NTG direct (see question 15), from TUG for $40 and from UKTUG for $30 (a manual is included). It is a useful resource for anyone to browse, not just for DOS users.

F TEX Systems

31 TEX/L^{A}TEX for different machines

We list here the free or shareware packages; see see question 33 for commercial packages.

Unix Instructions for retrieving the Unix TEX distribution via anonymous ftp are available in the document systems/

unix/unixtex.ftp

AIX TEX for the IBM RS6000 running AIX is available in

systems/unix/aix3.2

386/ix Executables for 386/ix are available in systems/

unix/386ix

Linux There are at least two fairly complete implementations of TEX to run on Linux. The Slackware distribution includes NTEX (available as systems/unix/linux/ ntex), which probably contains more TEX-related material than you would ever want. The more recent teTEX

(available as systems/unix/linux/teTeX) is based

on Karl Berry’s path-searching mechanisms, and is more compact than NTEX while still being pretty comprehensive.

PC The emTEX package for PCs running MS-DOS or OS/2 includes L^{A}TEX, BibTEX, previewers, and drivers, and is available in systems/msdos/emtex as a series of disk images. The package was written by Eberhard Mattes, and documentation is available in both German and English. Windows users, or those wishing to run current L^{A}TEX on machines with a 386 or better processor, should get the beta-test release of emTEX (systems/

msdos/emtex/betatest), and the RSX memory ex-
tender (systems/msdos/dpmigcc)

A second package, gTEX, runs under MS-DOS or Windows (and its users speak well of it). It is available from

systems/msdos/gtex

TUG (and some of the other user groups) offer all freelyavailable TEX software for the PC. A catalogue is available free from TUG (see question 14).

Mac OzTEX is a shareware version of TEX for the Macintosh. A dvi previewer and PostScript driver are also included. It should run on any Macintosh Plus, SE, II, or newer model, but will not work on a 128K or 512K Mac. It was written by Andrew Trevorrow, and is available in systems/mac/oztex, or on floppy disks from TUG (see question 14). UK TUG prepays the shareware fee, so that members of UK TUG may acquire the software without further payment. Questions about OzTEX may

be directed to oztex@midway.uchicago.edu

Another partly shareware program is CMacTEX (available as systems/mac/cmactex), put together by Tom Kiffe. This is much closer to the Unix TEX setup (it uses dvips, for instance).

VMS TEX for VMS is available as systems/vms/Alpha/ tex_axp_exe.zip (for Alpha-based machines) or

systems/vms/VAX/tex_vax_exe.zip (for VAX ma-

chines). Standard tape distribution is through DECUS (see question 27).

Atari TEX is available for the Atari ST in systems/atari

If anonymous ftp is not available to you, send a message containing the line “help” to atari@atari.archive.

umich.edu

Amiga Full implementations of TEX 3.1 (PasTEX) and - 2.7 are available in systems/amiga

You can also order a CD-ROM containing this and other Amiga software from Walnut Creek CDROM, telephone (+1) 510-947-5997.

TOPS-20 TEX was originally written on a DEC-10 under WAITS, and so was easily ported to TOPS-20. A distribution that runs on TOPS-20 is available via anonymous ftp from ftp.math.utah.edu (128.110.198.34)

in pub/tex/pub/web

32 TEX-friendly editors and shells

There are good TEX-writing environments and editors for most operating systems; some are described below, but this is only a personal selection:

Unix Try GNU emacs, and the AUCTEX mode (support/ auctex). This provides menu items and control sequences for common constructs, checks syntax, lays out markup nicely, lets you call TEX and drivers from within the editor, and everything else like this that you can think of. Complex, but very powerful.

VMS An lsedit mode for editing TEX source is available from TUG (see question 14) as TEXniques 1, VAX Language-Sensitive Editor, by Kent MacPherson (1985).

MS-DOS There are several choices:

• The 4AllTEX workbench (systems/msdos/ 4alltex) provides a very comprehensive environment written in 4DOS (you need to pay a shareware fee for this) which lets you access most TEX-related software in a friendly way. You can choose your own editor; something such as QEdit or Brief is suitable. This whole package is available in easyto-use form on CD-ROM from TEX user groups.

• TEXshell (systems/msdos/texshell) is a sim-

pler, easily-customisable environment, which can be used with the editor of your choice.

• Eddi4TEX (systems/msdos/e4t) is a speciallywritten TEX editor which features intelligent colouring, bracket matching, syntax checking, online help and the ability to call TEX programs from within the editor. It is highly customisable, and features a powerful macro language.

You can also use GNU emacs and AUCTEX under DOS.

Windows Your best public domain bet is probably to use MicroEmacs as an editor and control centre for TEX programs. The gTEX package (systems/msdos/gtex) comes with MicroEmacs ready to go, integrated with TEX, previewer, dvips and GhostScript.

On a PC with large enough memory, a version of GNU emacs, that will run under Windows, is available; thus you can also use AUCTEX under Windows.

Y&Y’s commercial (and high-quality) Windows previewer, dviwindo, can be used as a good TEX shell, calling programs such as TEX, drivers, and editors (Y&Y supply the public domain PE, and recommend the commercial Epsilon) from customisable menus (see question 33 for details of Y&Y).

Scientific Word is a wysiwyg editing program, strong on maths, which uses L^{A}TEX for output (see question 33 for contact address).

OS/2 Eddi4TEX works under OS/2; look also at systems/ os2/epmtex for a specific OS/2 shell.

Macintosh The commercial Textures provides an excellent integrated Macintosh environment with its own editor. More powerful still (as an editor) is the shareware Alpha

(systems/mac/support/alpha) which is extensible

enough to let you perform almost any TEX-related job. It works well with OzTEX.

Atari, Amiga and NeXT users also have nice environments. L^{A}TEX users who like ‘make’ can find help in, e.g., support/

latexmk

There is another set of shell programs to help you manipulate BibTEX databases.

33 Commercial TEX implementations

There are many commercial implementations of TEX. The first appeared not long after TEX itself appeared. Of the vendors, ArborText (formerly Textset) and Personal TEX are those who have survived longest (since the mid or early 80s).

What follows is probably an incomplete list. Naturally, no warranty or fitness for purpose is implied by the inclusion of any vendor in this list. The source of the information is given to provide some clues to its currency.

In general, a commercial implementation will come ‘complete’, that is, with suitable previewers and printer drivers. They normally also have extensive documentation (i.e., not just the TEXbook!) and some sort of support service. In some cases this is a toll free number (probably applicable only within the USA and or Canada), but others also have email, and normal telephone and fax support.

Unix; TEX Silicon Graphics Iris/Indigo, Solaris 2.1, IBM RS/6000, DEC/RISC-Ultrix, HP 9000. “Complete TEX packages. Ready to use, fully documented and supported.”

ArborText Inc 1000 Victors Way Suite 400 Ann Arbor MI 48108 USA Tel: (+1) 313-996-3566 Fax: (+1) 313-996-3573

Source: TUGboat, 1994, 15(1)

VAX/VMS; Convergent TEX Complete  system  for

VAX/VMS machines (a version for Alphas is in preparation); includes L^{A}TEX, multinational typesetting support, and Web.

Northlake Software, Inc. 812 SW Washington, Ste 1100 Portland, OR 97201 USA Tel: (+1) 503-228-3383 Fax: (+1) 503-228-5662

Email: rau@nls.com

Source: Email from Pat Rau, November 1994

PC; TurboTEX Runs on Windows, MS-DOS, OS/2, Unix and VAX/VMS; also TrueTEX for Windows 3.1, Window NT and Windows 95.

The Kinch Computer Co. 611 Mitchell Street Ithaca, New York 14850 USA Tel: (+1) 607-273-0222 Fax: (+1) 607-273-0484

Email: kinch@netcom.com

Source: AMS listing, November 1994

PC; TEX “Bitmap free TEX for Windows.”

Y&Y, Inc. 45 Walden Street Concord MA 01742 USA Tel: 800-742-4059 (within the USA) Tel: (+1) 508-371-3286 Fax: (+1) 508-371-2004

Source: AMS listing, November 1994

pcTEX Long-established: now has a Windows implementation.

Personal TEX Inc 12 Madrona Street Mill Valley, CA 94941 USA Fax: (+1) 415-388-8865

Email: pti@crl.com

Source: TUGboat, 1992, 13(2)

PC; VTEX Also “Bitmap-free”.

MicroPress Inc 68-30 Harrow Street Forest Hills, NY 11375 USA Tel: (+1) 718-575-1816 Fax: (+1) 718-575-8038 Source: Resource Directory supplement to TUGboat 1991, 12(12)

PC; microTEX MicroTEX and TEX tools.

Micro Programs, Inc. 251 Jackson Ave. Syosset, NY 11791 USA Tel: (+1) 516-921-1351

                                                     Email: sales@microprograms.com

Source: AMS listing, November 1994

PC; Scientific Word Scientific Word and Scientific Workplace offer a mechanism for near-wysiwyg input of L^{A}TEX documents; they ship with TrueTEX from Kinch (see above). Queries within the UK should be addressed to Scientific Word Ltd., others should be addressed directly to the publisher, TCI.

Dr Christopher Mabb Scientific Word Ltd. 98 Pont Adam Ruabon Wrexham Clwyd, LL14 6EF UK Tel: 01978 823088 (within the UK) Fax: 01978 823066 (within the UK)

                                                     Email: christopher@sciword.demon.
                                                     co.uk

TCI Software Research Inc.

1190 Foster Road

Las Cruces

New Mexico, 88001

USA

Tel: (+1) 505-522-4600 Fax: (+1) 505-522-0116

                                                     Email: info@tcisoft.com

Source: Mail from Christopher Mabb, November 1994

Macintosh; Textures “A TEX system ‘for the rest of us’ ”; also gives away a implementation and some font manipulation tools.

Blue Sky Research 534 SW Third Avenue Portland, Oregon 97204 USA Tel: 800-622-8398 (within the USA) Tel: (+1) 503-222-9571 Fax: (+1) 503-222-1643

                                                     Email: sales@bluesky.com

Source: TUGboat, 1994, 15(1)

AmigaTEX A full implementation for the Commodore Amiga, including full, on-screen and printing support for all PostScript graphics and fonts, IFF raster graphics, automatic font generation, and all of the standard macros and utilities.

Radical Eye Software PO Box 2081 Stanford, CA 94309 USA

Source: Mail from Tom Rokicki, November 1994

G DVI Drivers and Previewers

34 DVI to PostScript conversion programs

The best public domain dvi to PostScript conversion program which runs under many operating systems is Tom Rokicki’s dvips. dvips is written in C and ports easily to other operating systems; it is available as dviware/dvips

VMS versions are available through the DECUS library (see question 27), and also from CTAN: systems/

vms/Alpha/dvips555_axp_exe.zip (for Alpha-based
machines), systems/vms/VAX/dvips555_vax_exe.zip

(for VAX machines); support files are available in

systems/vms/dvips555_support.zip, and a set of fonts

for use with L^{A}TEX 2ε are available in systems/vms/

dvips_fontsupport.zip

A precompiled version for MS-DOS is available from

systems/msdos/dviware/dvips

Karl Berry’s version of dvips (called dvipsk) has a configure script and path searching code similar to that in his other programs (e.g., web2c); it is available from dviware/dvipsk

Another good portable program is dvitops by James Clark, which is also written in C and will compile under Unix, MS-DOS, VMS, and Primos; however, it does not support virtual fonts. It is available from dviware/dvitops

Macintosh users can use either the excellent drivers built into OzTEX or Textures, or a port of dvips in the CMacTEX package.

35 DVI drivers for HP LaserJet

The emTEX package (see question 31) contains a driver for the LaserJet, dvihplj.

Version 2.10 of the Beebe drivers supports the LaserJet. These drivers will compile under Unix, VMS, and on the Atari ST and DEC-20’s, and are available from dviware/beebe

Karl Berry’s dviljk, which has the same path-searching code as his dvipsk (see question 34), is available in dviware/

dviljk

36 DVI previewers

EmTEX and gTEX for the PC, and OzTEX for the Macintosh, all come with previewers that can be used on those platforms. There is a good OS/2 Presentation Manager previewer in emTEX, and a public domain Windows previewer (dviware/ dviwin). Commercial PC TEX packages (see question 33) have good DOS and Windows previewers.

Previewers available for other operating systems include:

xdvi The most widely used previewer for the X Window System (and hence almost any Unix or modern VMS workstation); available in dviware/xdvi Karl Berry’s version of xdvi, called xdvik, has features analogous to his dvipsk (see question 34); it is available

                                                         in dviware/xdvik

dvipage For SunView on (old enough) Sun workstations. This was published in volume 15 of comp.sources.unix

                                                         and is archived in dviware/dvipage

xtex An older previewer for the X Window System; available

                                                         in dviware/seetex

dviapollo For Apollo Domain workstations; available in

                                                         dviware/dviapollo

dvidis For (old enough, VMS) VAXstations running VWS;

                                                         available in dviware/dvidis

dvitovdu for Tektronix 4010-compatible and other terminals under Unix and VMS; available as dviware/dvitovdu

dvi2tty A dvi to ASCII conversion program, for normal terminals; available as dviware/dvi2tty

texsgi For SGI under Irix; both a binary and source are available, but be sure to get the fonts as well. Available as

                                                         dviware/texsgi

H Support Packages for TEX

37 Fig, a TEX-friendly drawing package

(X)Fig is a menu driven tool that allows you to draw objects on the screen of an X workstation. transfig is a set of tools which translate the code fig produces to other graphics languages including PostScript and the L^{A}TEX picture environment. They are

                                                     available in graphics/xfig and graphics/transfig

Fig is supported by Micah Beck (beck@cs.cornell.edu) and transfig is maintained by Brian Smith (bvsmith@lbl. gov). Another tool for fig conversion is fig2mf which generates code from fig input. It is available in graphics/

                                                     fig2mf

38 TEXCAD, a drawing package for L^{A}TEX

TEXCAD is a program for the PC which enables the user to draw diagrams on screen using a mouse or arrow keys, with an on-screen menu of available picture-elements. Its output is code for the L^{A}TEX picture environment. Optionally, it can be set to include lines at all angles using the emTEX driver-family \specials (see question 25). TEXCAD is part of the emTEX distribution.

A Linux port of the program, systems/unix/linux/ xtexcad-2.1.tar.z, is reported also to run on other Unix operating systems.

39 Spelling checkers for work with TEX

For Unix, ispell is probably the program of choice. It is available in support/ispell; beware of any version with a number 4.x — such versions represent a divergent version of the source which lacks many useful facilities of the 3.x series.

For MS-DOS, there are several programs. amspell can be called from within an editor (available as support/amspell). jspell is an extended version of ispell (available as support/

                                                     jspell).

For the Macintosh, Excalibur is the program of choice. It will run in native mode on both sorts of Macintosh,

and is available as systems/mac/support/excalibur/

Excalibur-20-sea.hqx (there are other dictionaries in the same directory).

For VMS, a spell checker can be found in support/

vmspell

40 The VORTEX package

VORTEX (available in support/vortex) is a package of programs written at the University of California at Berkeley, and was described by Michael A. Harrison in “News from the VORTEX project” in TUGboat 10(1), pp. 11–14, 1989. It includes several nice previewers and some emacs modes for TEX and BibTEX. The VORTEX distribution is not maintained, and now looks distinctly long in the tooth (it was never upgraded to TEX version 3).

I Literate programming

41 What is Literate Programming?

Literate programming is the combination of documentation and source together in a fashion suited for reading by human beings. In general, literate programs combine source and documentation in a single file. Literate programming tools then parse the file to produce either readable documentation or compilable source. The WEB style of literate programming was created by D. E. Knuth during the development of his TEX typesetting software.

Discussion of literate programming is conducted in the

newsgroup comp.programming.literate, which is gate-

wayed to the mailing list litprog@shsu.edu (see question 19 for details). The literate programming FAQ is stored as help/

LitProg-FAQ

42 WEB for C, FORTRAN, and other languages

TEX is written in the programming language WEB; WEB is a tool to implement the concept of “literate programming”.

CWEB, a WEB for C programs, written by Silvio Levy, is

available as web/c_cpp/cweb

Spidery WEB supports many languages including Ada, awk, and C. It was written by Norman Ramsey and, while not in the public domain, is usable free. It is available in web/

spiderweb

FWEB is a version for Fortran, Ratfor, and C written by John Krommes. It is available in web/fweb

SchemeWEB is a Unix filter that translates SchemeWEB into L^{A}TEX source or Scheme source. It was written by John Ramsdell and is available in web/schemeweb

APLWEB is a version of WEB for APL and is available in

web/apl/aplweb

FunnelWeb is a version of WEB that is language independent. It is available in web/funnelweb

Other language independent versions of WEB are nuweb (which is written in ANSI C), available in web/nuweb, and noweb, available in web/noweb

J Format conversions

43 Conversion between TEX/L^{A}TEX and others

troff troff-to-latex (available as support/troff-to-latex), written by Kamal Al-Yahya at Stanford University (California, USA), assists in the translation of a troff document into L^{A}TEX format. It recognises most -ms and -man macros, plus most eqn and some tbl preprocessor commands. Anything fancier needs to be done by hand. Two style files are provided. There is also a man page (which converts very well to L^{A}TEX...). The program is copyrighted but free. An enhanced version of this program, tr2latex, is available in support/tr2latex

The DECUS TEX distribution (see question 27) also contains a program which converts troff to TEX.

Scribe Mark James (jamesm@dialogic.com) has a copy of scribe2latex he has been unable to test but which he will let anyone interested have. The program was written by Van Jacobson of Lawrence Berkeley Laboratory.

WordPerfect wp2latex (available as support/wp2latex) is a PC program written in Turbo Pascal by R. C. Houtepen at the Eindhoven University in the Netherlands. It converts WordPerfect 5.0 documents to L^{A}TEX. Pascal source is included. Users find it “helpful” and “decent” in spite of some limitations. It gets high marks for handling font changes, but cannot make indices, tables of contents, margins or graphics, and can’t handle features new in WordPerfect version 5.1, in particular the equation formatter. The program is copyrighted but free.

Glenn Geers of the University of Sydney, Aus-

tralia (glenn@qed.physics.su.oz.au) is translating

wp2latex into C and adding some WordPerfect 5.1 features, in particular its equation handling. His work is in

the glenn subdirectory of support/wp2latex

PC-Write pcwritex.arc is available in support/

pcwritex

This is a print driver for PC-Write that “prints” a PC-Write V2.71 document to a TEX-compatible disk file. It was written by Peter Flynn at University College, Cork, Republic of Ireland.

runoff Peter Vanroose’s (vanroose@esat.kuleuven.ac.

be) conversion program is written in VMS Pascal. The sources and a VAX executable are available from

support/rnototex

refer/tib There are a few programs for converting bibliographic data between BibTEX and refer/tib formats.

They are in biblio/bibtex/utils/refer-tools

In spite of the directory name, it also contains a shell script to convert BibTEX to refer as well. The collection is not maintained.

RTF A program for converting Microsoft’s Rich Text Format to TEX is available in support/rtf2tex, which was written and is maintained by Robert Lupton (rhl@ astro.princeton.edu). There is also a convertor to L^{A}TEX by Erwin Wechtl, in support/rtf2latex Microsoft Word A rudimentary program for converting MS-

Word to L^{A}TEX is wd2latex, for MS-DOS (dviware/ wd2latex); a better idea, however, is to convert the document to RTF format and use one of the RTF converters mentioned above.

A group at Ohio State University (USA) is working on a common document format based on SGML, with the ambition that any format could be translated to or from this one. Framemaker provides “import filters” to aid translation from alien formats (presumably including TEX) to Framemaker’s own.

44 Conversion from TEX/L^{A}TEX to plain ASCII

The aim here is to emulate the Unix nroff, which formats text as best it can for the screen, from the same input as the Unix typesetting program troff.

Ralph Droms (droms@bucknell.edu) has a style file and a program that provide the L^{A}TEX equivalent of nroff, though it doesn’t do a good job with tables and mathematics. The software is available in support/txt;the original dvi2tty often does an acceptable job and is available in dviware/dvi2tty

Another possibility is to use screen.sty (available as

macros/latex209/contrib/misc/screen.sty). Use a

dvi2tty program of some kind; you might try dviware/ crudetype as well. Another possibility is to use the L^{A}TEX-to-ASCII conversion program, l2a (support/l2a), although this is really more of a de-TEXing program.

The canonical de-TEXing program is detex (support/ detex), which removes all comments and control sequences from its input before writing it to its output. Its original purpose was to prepare input for a dumb spelling checker.

45 Conversion from HTML or SGML to TEX

Nathan Torkington (Nathan.Torkington@vuw.ac.nz) de-

veloped html2latex from the HTML parser in NCSA’s Xmosaic package. The program takes an HTML file and generates a L^{A}TEX file from it. The conversion code is subject NCSA restrictions, but the whole source is available as support/

html2latex

Peter Flynn has developed a pcl program sgml2tex (for PCs). It is controlled by a configuration file, and an example configuration for translating HTML files is given. The program

is available in support/sgml2tex

46 Conversion to HTML

L^{A}TEX2HTML (support/latex2html) is a package (mostly of perl scripts) that breaks up a L^{A}TEX document into one or more components, and links them together so that they can be read over the World-Wide Web as an hypertext document.

L^{A}TEX2HTML defines a mapping between L^{A}TEX intradocument references and hyperlinks, and extends the L^{A}TEX referencing mechanisms to permit reference to other (possibly remote) documents and other internet resources. It translates L^{A}TEX accented and other characters (as best it can) to things that World-Wide Web browsers can display, and provides a mechanism whereby mathematics (and other things that browsers can’t deal with) may be translated to images that can be loaded in-line into the hypertext document.

K ��������

47 Getting �������� to do what you want

allows you to create your own fonts, and most TEX users will never need to use it. , unlike TEX, requires some customisation: each output device for which you will be generating fonts needs a mode associated with it. Modes are defined using the mode_def convention described on page 94 of The book (see question 17). You will need a file, which conventionally called local.mf, containing all the mode_defs you will be using. If local.mf doesn’t already exist, Karl Berry’s collection of modes, available as fonts/modes/modes.mf, is a good starting point (it can be used as a ‘local.mf’ without modification in a ‘big enough’ implementation of ). Lists of settings for various output devices are also published periodically in TUGboat (see question 14). Now create a plain base file using

                                                     inimf, plain.mf, and local.mf:
                                                        % inimf

This is METAFONT...

**plain  you type ‘plain’

(output)

*input local  you type this

(output)

*dump  you type this

Beginning to dump on file plain...

(output)

This will create a base file named plain.base (or something similar; for example, it will be PLAIN.BAS on MS-DOS systems) which should be moved to the directory containing the base files on your system (note that some systems have two or more such directories, one for each ‘size’ of used). Now you need to make sure loads this new base when it starts up. If loads the plain base by default on your system, then you’re ready to go. Under Unix (using the default web2c distribution^{8}) this does indeed happen, but we could for instance define a command mf which executes virmf &plain loading the plain base file.

The usual way to create a font with plain is to start it with the line

                                                       \mode=<mode name>;mag=<magnification>;
                                                          input <font file name>

in response to the ‘**’ prompt or on the command line. (If <mode name> is unknown or omitted, the mode defaults to ‘proof’ and will produce

                                                     an output file called <font file name>.2602gf) The

<magnification> is a floating point number or ‘magstep’ (magsteps are defined in The book and The TEXbook). If mag=<magnification> is omitted, then the default is 1 (magstep 0). For example, to generate cmr10 at 12pt for an epson printer you would type

8The command name is symbolically linked to virmf, and virmf loads command_name.base

mf \mode=epson; mag=magstep 1; input cmr10

Note that under Unix the \ and ; characters must usually be quoted or escaped, so this would typically look something like

mf '\mode=epson; mag=magstep 1; input cmr10'

If you don’t have inimf or need a special mode that isn’t in the base, you can put its commands in a file (e.g., ln03.mf) and invoke it on the fly with the \smode command. For example, to create cmr10.300gf for an LN03 printer, using the file

% This is ln03.mf as of 2/27/90
% mode_def courtesy of John Sauter
proofing:=0;
fontmaking:=1;
tracingtitles:=0;
pixels_per_inch:=300;
blacker:=0.65;
fillin:=-0.1;
o_correction:=.5;

(note the absence of the mode_def and enddef commands), you would type

mf \smode="ln03"; input cmr10

This technique isn’t one you should regularly use, but it may prove useful if you acquire a new printer and want to experiment with parameters, or for some other reason are regularly editing the parameters you’re using. Once you’ve settled on an appropriate set of parameters, you should use them to rebuild the base file that you use.

A summary of the above written by Geoffrey Tobin, and tips about common pitfalls in using , is available

as documentation/metafont_for_beginners.tex

48 Which font files should be kept

produces from its run three files, a metrics (tfm) file, a generic font (gf) file, and a log file; all of these files have the same base name as does the input (e.g., if the input file was

cmr10.mf, the outputs will be cmr10.tfm, cmr10.nnngf^{9} and
cmr10.log).

To operate TEX using the font, you need a tfm file, so you need to keep that. However, you are likely to want generate the same font at more than one magnification, and each time you do so you’ll (incidentally) generate another tfm file; these files are all the same, and you only therefore need to retain one of them.

To produce preview or printed output, the dvi processor will need a font raster file; this is what the gf file provides. However, while there used (once upon a time) to be dvi processors that could use gf files, modern processors use packed raster (pk) files. Therefore, you need to generate a pk file from the gf file; the program gftopk does this for you, and once you’ve done that you may throw the gf file away.

The log file should never need to be used, unless there was some sort of problem in the run, and need not be ordinarily kept.

49 Getting bitmaps from the archives

Most people these days start using TEX with a 300 dots-per-inch (dpi) laser printer, and Computer Modern bitmap fonts for this resolution are supplied with most TEX packages. There are also two such sets available on CTAN: fonts/cm/pk/ pk300.zip (for write-black printer engines) and fonts/cm/ pk/pk300w.zip (for write-white engines). However, some users want to send their work to high quality typesetting machines (typically with a resolution of 1270 dpi or greater); it is also becoming more common to use a 600 dpi laser printer. Why don’t the archives or suppliers provide bitmap fonts at these sizes? There are two reasons:

1. When a bitmap font is created with , it needs to know the characteristics of the device; who knows what 600 or 1270 dpi device you have? (Of course, this objection applies equally well to 300 dpi printers.)

2. Bitmap fonts get big at high resolutions. Who knows what fonts at what sizes you need?

It would be possible to provide some set of 1270 dpi bitmap fonts in the archives, but it would take a lot of space, and might not be right for you.

So what to do? You can build the fonts you need yourself with : this isn’t at all hard, and some drivers help you (dvips, and the emTEX drivers) construct the commands. You might need to look at Karl Berry’s collection

of  modes (fonts/modes/modes.mf). Alterna-

tively, if it is a PostScript device you have, consider using the fonts in Type 1 font format. You can buy all the Computer Modern fonts in PostScript form from Blue Sky Research or Y&Y (see question 33 for addresses), or you can use Basil Malyshev’s public domain versions in fonts/cm/ps-type1 (the Paradissa collection is complete, but is gradually being replaced by the better BaKoMa collection).

L PostScript and TEX

50 Using PostScript fonts with TEX

Before you start: in order to use PostScript fonts, TEX needs metric (called tfm) files. Several sets of metrics are available from the archives; for mechanisms for generating new ones, see question 52. You also need the fonts themselves; PostScript printers come with a set of fonts built in, but to extend your repertoire you almost invariably need to buy from one of the many commercial font vendors (see, for example, question 54).

If you use L^{A}TEX 2ε, the best way to get PostScript fonts into your document is to use the PSNFSS package maintained by Sebastian Rahtz and Alan Jeffrey (available in macros/latex/ packages/psnfss); it’s supported by the L^{A}TEX3 project team, so bug reports can and should be submitted. PSNFSS gives you a set of packages for changing the default roman, sansserif and typewriter fonts; e.g., times.sty will set up Times Roman, Helvetica and Courier in place of Computer Modern, while avant.sty just changes the sans-serif family to Avant-Garde. To go with these packages, you will need the font metric files (watch out for encoding problems! see question 52) and font description (.fd) files for each font family you want to use. These can be obtained from fonts/metrics, arranged by vendor (e.g., Adobe, Monotype, etc.). For convenience, metrics for the common ‘35’ PostScript fonts found in most printers are provided with PSNFSS, packaged as macros/latex/

9Note that the file name may be transmuted by such operating systems as MS-DOS, which don’t permit long file names

packages/psnfss/lw35nfss.zip

For older versions of L^{A}TEX there are various schemes, of which the simplest to use is probably the PSL^{A}TEX macros distributed with dvips.

For plain TEX, you load whatever fonts you like; if the encoding of the fonts is not the same as Computer Modern it will be up to you to redefine various macros and accents, or you can use the font re-encoding mechanisms available in many drivers and in ps2pk and afm2tfm.

Victor Eijkhout’s sophisticated Lollipop package (macros/ lollipop) supports declaration of font families and styles in a similar way to L^{A}TEX’s NFSS, and so is easy to use with PostScript fonts.

Some common problems encountered are discussed elsewhere (see question 53).

51 Previewing files using PostScript fonts

Most TEX previewers only display bitmap pk fonts. If you want to preview documents using PostScript fonts, you have three choices:

1. Convert the .dvi file to PostScript and use a PostScript previewer. Some modern Unix X implementations have this built in (as does NeXT-step); (X11) Unix, Windows, OS/2, and MS-DOS users can use the free GhostScript (support/ghostscript), a complete level 2 implementation.

2. Under Windows on a PC, or on a Macintosh, let Adobe Type Manager display the fonts. Textures (Macintosh) works like this, and under Windows you can use Y&Y’s dviwindo for bitmap-free previewing. (See question 33 for details of these suppliers.)

3. If you have the PostScript fonts in Type 1 format, use ps2pk (fonts/utilities/ps2pk) or gsftopk (designed for use with the ghostscript fonts; fonts/utilities/ gsftopk) to make pk bitmap fonts which your previewer will understand. This can produce excellent results, also suitable for printing with on non-PostScript devices. Check the legalities of this if you have purchased the fonts. The very commonest PostScript fonts such as Times and Courier come in Type 1 format on disk with Adobe Type Manager (often bundled with Windows, and part of OS/2).

52 TEX font metric files for PostScript fonts

Font vendors such as Adobe supply metric files for each font, in afm (Adobe Font Metric) form; these can be converted to tfm (TEX Font Metric) form. The CTAN archives have prebuilt metrics which will be more than enough for many people (fonts/ metrics; beware — this directory is at the root of a huge tree), but you may need to do the conversion yourself if you have special needs or acquire a new font. One important question is the encoding of (Latin character) fonts; while we all more or less agree about the position of about 96 characters in fonts (the basic ASCII set), the rest of the (typically) 256 vary. The most obvious problems are with floating accents and special characters such as the ‘pounds sterling’ sign. There are three ways of dealing with this: either you change the TEX macros which reference the characters (not much fun, and error-prone); or you change the encoding of the font (easier than you might think); or you use virtual fonts (see question 24) to pretend to TEX that the encoding is the same as it is used to. If you use L^{A}TEX 2ε, it allows for changing the encoding in TEX; read the L^{A}TEX Companion (see question 17) for more details. In practice, if you do much non-English (but Latin script) typesetting, you are strongly recommended to use the fontenc package with option ‘T1’ to select T1 (also known as ‘Cork’ or ‘DC’) encoding. The PSNFSS package (see question 50) provides support for PostScript (using virtual fonts) in this encoding.

Alan Jeffrey’s fontinst package (fonts/utilities/ fontinst) is an afm to tfm converter written in TEX; it is used to generate the files used by L^{A}TEX 2ε’s PSNFSS package to support use of PostScript fonts. It is a sophisticated package, not for the faint-hearted, but is powerful enough to cope with most needs. Much of its power relies on the use of virtual fonts (see question 24).

For slightly simpler problems, Rokicki’s afm2tfm, distributed with dvips (dviware/dvips), is fast and efficient; note that the metrics and styles that come with dvips are not currently L^{A}TEX 2ε compatible.

For the Macintosh, there is a program called EdMetrics which does the job (and more). It comes with the Textures distribution, but is in fact free software, available as systems/

mac/textures/utilities/EdMetrics.sea.hqx

MS-DOS users can buy (see question 33) Y&Y’s Font Manipulation Tools package which includes a powerful afmtotfm program among many other goodies.

53 Problems using PostScript fonts

For the typical L^{A}TEX user trying to use the PSNFSS (see question 50) package, three questions often arise. First, you have to declare to the dvi driver that you are using PostScript fonts; in the case of dvips, this means adding lines to the psfonts.map file. Otherwise, dvips will try to find pk files. If the font isn’t built into the printer, you have to acquire it (in many cases this means buying it from a commercial supplier!). You then have to instruct the driver to download it with each job (the mechanism depends on your driver). So it’s no good just installing the metrics for Optima and expecting it to work. You have to pay hard cash for the font itself, which will come (for Unix and MS-DOS users) in pfb (Printer Font Binary) form.

Second, you cannot expect your previewer to suddenly start displaying PostScript fonts; most of them only know about PK bitmap fonts such as Computer Modern. You can create these from PostScript fonts if you have bought the pfb file by using ps2pk (fonts/utilities/ps2pk); this would also let you use the fonts with non-PostScript device drivers such as the emTEX ones. It is your responsibility to make sure you are not breaking the licence restrictions on your font if you bought it.

Third, the stretch and shrink between words is a function of the font metric; it is not specified in afm files, so different converters choose different values. The PostScript metrics that come with PSNFSS do quite tight setting, which may produce a lot of hyphenation or overfull boxes if you are not careful. If you are having trouble, Alan Jeffrey recommends the following settings in a package file:

\tolerance800
\emergencystretch2em
\doublehyphendemerits5000
\hfuzz0pt
\leftskip0pt minus 1pt
\rightskip0pt minus 1pt

A future release of these fonts will implement looser typesetting by default, so these settings should become unnecessary.

54 Choice of scalable outline fonts

If you are interested in text alone, you can use any of over 20,000 fonts(!) in Adobe Type 1 format (called ‘PostScript fonts’ in the TEX world and ‘ATM fonts’ in the DTP world), or any of several hundred fonts in TrueType format. That is, provided of course, that your previewer and printer driver support scalable outline fonts.

TEX itself only cares about metrics, not the actual character programs. You just need to create a TEX metric file .tfm using some tool such as afm2tfm, afmtotfm (from Y&Y, see question 33) or fontinst. For the previewer or printer driver you need the actual outline font files themselves (pfa for Display PostScript, pfb for ATM on IBM PC, Mac outline font files on Macintosh).

If you also need mathematics, then you are severely limited by the demands that TEX makes of maths fonts (for details, see the paper by B.K.P. Horn in the Proceedings of the 1993 TUG conference at Aston University, published in TUGboat, 1994, 14(3)). For maths, then, there are relatively few choices:

Computer Modern (75 fonts — optical scaling) Donald E. Knuth Note that CM is available in scalable outline form. There are commercial as well as public domain versions, and there are both Adobe Type 1 and TrueType versions. Some of these are ‘commercial grade,’ with full handtuned hinting, some render very poorly, while others are merely incompatible with Adobe Type Manager (ATM).

Lucida Bright with Lucida New Math (25 fonts) Chuck Bigelow and Kris Holmes Lucida is a family of related fonts including seriffed, sans serif, sans serif fixed width, calligraphic, blackletter, fax, Kris Holmes’ connected handwriting font, etc; they’re not as ‘spindly’ as Computer Modern, with a large x-height, and include a larger set of maths symbols, operators, relations and delimiters than CM (over 800 instead of 384: among others, it also includes the AMS msam and msbm symbol sets). The planned ‘Lucida Bright Expert’ (14 fonts) adds seriffed fixed width, another handwriting font, smallcaps, bold maths, upright ‘maths italic’, etc., to the set The distribution includes support for use with plain TEX and L^{A}TEX 2.09. Support under L^{A}TEX 2ε is provided in PSNFSS (see question 50) thanks to Sebastian Rahtz.

MathTime 1.1 (3 fonts) TEXplorators (Michael Spivak) The set contains maths italic, symbol, and extension fonts, designed to work well with Times-Roman. These are typically used with Times, Helvetica and Courier (which are resident on many printers, and which are supplied with some PC versions). In addition you may want to complement this basic set with Adobe’s Times Smallcap, and perhaps the set of Adobe ‘Math Pi’ fonts, which include blackboard bold, blackletter, and script faces. The distribution includes support for use with plain TEX and L^{A}TEX 2.09 (including code to link in Adobe Math Pi 2 and Math Pi 6). Support under L^{A}TEX 2ε is provided in PSNFSS (see question 50) thanks to Sebastian Rahtz.

Adobe Lucida, LucidaSans and LucidaMath (12 fonts) Lucida and LucidaMath are generally considered to be a bit heavy. The three maths fonts contain only the glyphs in the CM maths italic, symbol, and extension fonts. Support for using LucidaMath with TEX is not very good; you will need to do some work reencoding fonts etc. (In some sense this set is the ancestor of the LucidaBright plus LucidaNewMath font set.)

Concrete, the AMS maths fonts etc. Donald E. Knuth and the AMS. These are sometimes mentioned as alternatives to CM, but they are really adjuncts, in that you need to use at least the basic CM maths fonts with them.

Proprietary fonts Various sources.

Since having a high quality font set in scalable outline form that works with TEX can give a publisher a real competitive advantage, there are some publishers that have paid (a lot) to have such font sets made for them. Unfortunately, these sets are not available on the open market, despite the likelihood that they’re more complete than those that are.

Mathptm (4 fonts) Alan Jeffrey.

This set contains maths italic, symbol, extension, and roman virtual fonts, built from Adobe Times, Symbol, Zapf Chancery, and the Computer Modern fonts. The Mathptm fonts are free, and the resulting PostScript files can be freely exchanged. Contains most of the CM math symbols. Support under L^{A}TEX 2ε in PSNFSS (see question 50) thanks to Alan Jeffrey and Sebastian Rahtz.

All of the first three font sets are available in formats suitable for IBM PC/Windows, Macintosh and Unix/NeXT from Y&Y and from Blue Sky Research (see question 33 for details). The MathTime fonts are also available from:

TEXplorators 1572 West Gray #377 Houston TX 77019 USA

The very limited selection of maths font sets is a direct result of the fact that a maths font has to be explicitly designed for use with TEX and as a result it is likely to lose some of its appeal in other markets. Furthermore, the TEX market for commercial fonts is minute (in comparison, for example, to Microsoft True-Type font pack #1, which sold something like 10 million copies in a few weeks after release of Windows 3.1!).

Text fonts in Type 1 format are available from many vendors including Adobe, Monotype, Bitstream. Avoid cheap ripoffs: not only are you rewarding unethical behaviour, destroying the cottage industry of innovative type design, but you are also very likely to get junk. The fonts may not render well (or at all under ATM), may not have the ‘standard’ complement of 228 glyphs, or may not include metric files (needed to make TFM files). Also, avoid TrueType fonts from all but the major vendors. TrueType fonts are an order of magnitude harder to ‘hint’ properly than Type 1 fonts and hence TrueType fonts from places other than Microsoft and Apple may be suspect. In any case you may find other problems with TrueType fonts such as service bureaux not accepting jobs calling for them.

55 Including a PostScript figure in L^{A}TEX

L^{A}TEX 2ε (see question 95) has a standard package for graphics inclusion, rotation, colour, and other driver-related features. The package is documented in the second edition of the Lamport’s L^{A}TEX book (see question 17), and is available in macros/

latex/packages/graphics

If you don’t use L^{A}TEX 2ε, perhaps the best method is to use the psfig macros written by Trevor Darrell, available in

graphics/psfig

You will also need a dvi to PostScript conversion program that supports the \specials. The drivers mentioned in question 34 do, and come with a version of psfig ready to use with them. The psfig macros work best with Encapsulated PostScript Files (EPS). In particular, psfig will need the file to have a BoundingBox (see Appendix H of the PostScript Language Reference Manual). If you don’t have an EPS file, life can be difficult.

One point to note about including PostScript figures is that they are not part of the dvi file, but are only included when you use a dvi to PostScript conversion program. As a result, most dvi previewers will simply show the blank space TEX has reserved for your figure, not the figure itself.

Anil K. Goel has written a long document describing in detail how to include figures, pictures, and images in L^{A}TEX documents, which is available as info/figsinltx.ps

M Special sorts of typesetting

56 Double-spaced documents in L^{A}TEX

Are you producing a thesis, and trying to obey regulations that were drafted in the typewriter era? Or are you producing copy for a journal that insists on double spacing for the submitted articles?

L^{A}TEX is a typesetting system, so the appropriate design conventions are for “real books”. If your requirement is from thesis regulations, find whoever is responsible for the regulations, and try to get the wording changed to cater for typeset theses (e.g., to say “if using a typesetting system, aim to make your thesis look like a well-designed book”). (If your requirement is from a journal, you’re probably even less likely to be able to get the rules changed, of course.)

If you fail to convince your officials, or want some inter-line space for copy-editing:

   • Try changing \baselinestretch: \renewcommand
     {\baselinestretch}{1.2} may be enough to give

officials the impression you’ve kept to their regulations. Don’t try changing \baselineskip: its value is reset at any size-changing command.

• Alternatively, use a line-spacing package. Options available are:

– for simple double spacing, macros/latex209/

                                                              contrib/misc/doublespace.sty, and

– for greater flexibility, macros/latex/contrib/

                                                              other/misc/setspace.sty, which has been up-

graded for L^{A}TEX 2ε.

57 Formatting a thesis in L^{A}TEX

Thesis styles are usually very specific to your University, so it’s usually not profitable to ask around for a package outside your own University. Since many Universities (in their eccentric way) still require double-spacing, you may care to refer to question 56. If you want to write your own, a good place to start is the University of California style (available as

                                                     macros/latex209/contrib/ucthesis), but it’s not worth

going to a lot of trouble. (If officials won’t allow standard typographic conventions, you won’t be able to produce an æsthetically pleasing document anyway!)

58 Including a file in verbatim in L^{A}TEX

A good way is to use Rainer Schöpf’s verbatim.sty, which provides the command \verbatiminput that takes a file name as argument. This package is available as part of macros/

                                                     latex/packages/tools

Another way is to use the alltt environment, which requires the package available in macros/latex209/contrib/

                                                     misc/alltt.sty

59 Generating an index in TEX/L^{A}TEX

Making an index is not trivial; what to index, and how index it, is difficult to decide, and uniform implementation is difficult to achieve. You will need to mark all items to be indexed in your text (typically with \index commands).

It is not practical to sort a large index within TEX, so a postprocessing program is used to sort the output of one TEX run, to be included into the document at the next run. The following programs are available:

makeindex for L^{A}TEX under Unix (but runs under other OSs without changes). Available in indexing/makeindex; a version for the Macintosh is available as systems/

                                                         mac/macmakeindex.sit, and ones for MS-DOS are

part of the emTEX and gTEX distributions (the emTEX version also runs under OS/2).

The Makeindex documentation is a good source of information on how to create your own index. Makeindex can be used with some TEX macro packages other than L^{A}TEX, such as Eplain.

idxtex for L^{A}TEX under VMS. Available (together with a glossary-maker called glotex) in indexing/

                                                         glo+idxtex

texindex A witty little shell/sed-script-based utility for L^{A}TEX under Unix. Available from support/texindex

There are other programs called texindex, notably one that comes with the Texinfo distribution (see question 11).

60 Using BibTEX with plain TEX

The file macros/eplain/btxmac.tex contains macros and

documentation for using BibTEX with plain TEX, either directly or with Eplain (see question 9). See question 20 for more information about BibTEX itself.

61 Typesetting music in TEX

A powerful package which allows the typesetting of orchestral and polyphonic music is MusicTEX, written by Daniel

Taupin (taupin@rsovax.lps.u-psud.fr). It is available in
macros/musictex

Digital music fans can typeset notation for their efforts by using midi2tex, which translates MIDI data files into MusicTEX source code. It is available from support/midi2tex

A rather simpler notation than MusicTEX is supported by abc2mtex; this is a package designed to notate tunes stored in an ASCII format (abc notation). It was designed primarily for folk and traditional tunes of Western European origin (such as Irish, English and Scottish) which can be written on one stave in standard classical notation. However, it should be extendable to many other types of music. It is available from support/

abc2mtex

There is a mailing list for discussion of typesetting music in TEX. To subscribe, send mail to mutex-request@stolaf. edu containing the word ‘subscribe’ in the body.

62 Drawing Feynman diagrams in L^{A}TEX

Michael Levine’s macro package for drawing Feynman diagrams in L^{A}TEX is available in macros/latex209/contrib/

feynman (see also macros/latex/contrib/supported/
feynmf).

N How do I do X in TEX or L^{A}TEX

63 Proof environment

It is not possible to make a proof environment which automatically includes an ‘end-of-proof’ symbol. Some proofs end in displayed maths; others do not. If the input file contains ...\] \end{proof} then L^{A}TEX finishes off the displayed maths and gets ready for a new line before it reads any instructions connected with ending the proof. But traditionally the end-of-proof sign goes in the display, not on a new line. So you just have to put it in by hand in every proof.

64 Symbols for the number sets

It is a good idea to have commands such as \R for the real numbers and other standard number sets. Traditionally these were typeset in bold. Because mathematicians usually do not have access to bold chalk, they invented the special symbols that are now often used for \R, \C, etc. These symbols are known as “blackboard bold”. Before insisting on using them, consider whether going back to the old system of ordinary bold might not be acceptable (it is certainly simpler).

A set of blackboard bold capitals is available in the AMS fonts “msam” (e.g., “msam10” for 10pt) and “msbm”. The fonts have a large number of mathematical symbols to supplement the ones in the standard TEX distribution. The fonts are

                                                     available in fonts/ams/amsfonts/sources/symbols

Two files which load the fonts and define the symbols are provided, and both work with either TEX or L^{A}TEX. Questions or suggestions regarding these fonts should be directed to tech-

                                                     support@math.ams.org.

Another complete set of blackboard bold fonts, the bbold family, is available in (in fonts/bbold). This set has the interesting property of offering blackboard bold forms of lower-case letters, something rather rarely seen on actual blackboards.

The “lazy person’s” blackboard bold macros:

                                                       \newcommand{\R}{{\sf R\hspace*{-0.9ex}%
                                                         \rule{0.15ex}{1.5ex}\hspace*{0.9ex}}}
                                                       \newcommand{\N}{{\sf N\hspace*{-1.0ex}%
                                                         \rule{0.15ex}{1.3ex}\hspace*{1.0ex}}}
                                                       \newcommand{\Q}{{\sf Q\hspace*{-1.1ex}%
                                                         \rule{0.15ex}{1.5ex}\hspace*{1.1ex}}}
                                                       \newcommand{\C}{{\sf C\hspace*{-0.9ex}%
                                                         \rule{0.15ex}{1.3ex}\hspace*{0.9ex}}}

work well at normal size if the surrounding text is cmr10. However, they are not part of a proper maths font, and so do not work in sub- and superscripts. Moreover, the size and position of the vertical bar is affected by the font of the surrounding text.

65 Roman theorems

If you want to take advantage of the powerful \newtheorem command without the constraint that the contents of the theorem is in a sloped font (for example, to use it to create remarks, examples, proofs, ...) then you can use the style file

                                                     theorem.sty (part of macros/latex/packages/tools).

Alternatively, the following sets up an environment remark whose content is in roman.

                                                       \newtheorem{preremark}{Remark}
                                                       \newenvironment{remark}%
                                                         {\begin{preremark}\rm}{\end{preremark}}

This will not work if you are using NFSS (see question 94) outside of L^{A}TEX 2ε (see question 95), because the command \rm behaves differently there.

66 Labels on lists

If you want your top-level enumerates to be labelled ‘I/’, ‘II/’, ..., then give these commands:

                                                        \renewcommand{\theenumi}{\Roman{enumi}}
                                                        \renewcommand{\labelenumi}{\theenumi/}

The possible styles of numbering are given in Section 6.3 of Lamport’s book (see question 17). If you are lazy and just change \labelenumi in one go then your cross-references will be wrong.

For lower level enumerates, replace enumi by enumii, enumiii or enumiv, according to the level. If your label is much larger than the default, you should also change

\leftmargini, \leftmarginii, etc.

If you’re running L^{A}TEX 2ε, the package enumerate.sty

(part of macros/latex/packages/tools) offers simi-

lar facilities. With enumerate.sty, the example above would be achieved simply by starting the enumeration

\begin{enumerate}[I/].

67 Unnumbered sections in the Table of Contents

The easiest way to get headings of funny ‘sections’ such as prefaces in the table of contents is to use the counter secnumdepth described in Appendix C of the L^{A}TEX manual. For example:

\setcounter{secnumdepth}{-1}
\chapter{Preface}

Of course, you have to set secnumdepth back to its usual value (which is 2 in the standard styles) before you do any ‘section’ which you want to be numbered.

This is why it works. \chapter without the star does

1. put something in the .toc file;

2. if secnumdepth ≥ 0, increase the counter for the chapter and write it out.

3. write the chapter title.

Other sectioning commands are similar, but with other values used in the test.

68 Style of section headings

Suppose that the editor of your favourite journal has specified that section headings must be centred, in small capitals, and subsection headings ragged right in italic, but that you don’t want to get involved in the sort of programming described in The L^{A}TEX Companion (see question 17; the programming itself is discussed in question 85). The following hack will probably satisfy your editor. Define yourself new commands

\newcommand{\ssection}[1]{%
\section[#1]{\centering\sc #1}}
\newcommand{\ssubsection}[1]{%
\subsection[#1]{\raggedright\it #1}}

and then use \ssection and \ssubsection in place of \section and \subsection. This isn’t perfect: section numbers remain in bold, and starred forms need a separate redefinition. Also, this will not work if you are using NFSS (see question 94) outside of L^{A}TEX 2ε (see question 95), because the fontchanging commands behave differently there.

69 Indent after section headings

L^{A}TEX implements a style that doesn’t indent the first paragraph after a section heading. There are coherent reasons for this, but not everyone likes it. The package indentfirst.sty (part

of macros/latex/packages/tools) suppresses the mecha-

nism, so that the first paragraph is indented.

70 Changing the margins in L^{A}TEX

Don’t do it. Learn some L^{A}TEX, produce some documents, and then ask again.

You can never change the margins of a document by software, because these depend on the actual size of the paper, which may be different from what your software thinks. (For example, standard US paper is a different width from A4 paper.) What you can change are the distances from the apparent top and left edges of the paper, and the width and height of the text. Changing the last two requires more skill than you might expect. The height should bear a certain relationship to \baselineskip. And the width should not be more than 75 characters. Lamport’s warning in his section on ‘Customizing the Style’ really must be taken seriously. One-inch margins on A4 paper are fine for 10- or 12-pitch typewriters, but not for 10pt type (or even 11pt or 12pt) because so many characters per line will irritate the reader.

71 Insisting on changing the margins in L^{A}TEX

This answer first helps you change the margins throughout a document, then tells you how to change the margins in a portion of the document.

Perhaps the easiest way to get more out of a page in L^{A}TEX

is to get macros/latex209/contrib/misc/fullpage.sty,

which sets the margins of the page identical to those of plain TEX, i.e., 1-inch margins at all four sides of the paper. It also contains an adjustment for A4 paper.

   Somewhat more flexible is macros/latex/contrib/

other/misc/vmargin.sty, which has a canned set of paper sizes (a superset of that provided in L^{A}TEX 2ε), provision for custom paper, margin adjustments and provision for two-sided printing.

Here is a brief explanation of L^{A}TEX’s page parameters. For details, see section C.5.3 of the L^{A}TEX manual (pp. 181– 182). The origin in dvi coordinates is one inch from the top of the paper and one inch from the left side. This explains the “one inch less than” terminology used in the L^{A}TEX manual. In dvi coordinates, positive horizontal measurements extend right across the page, and positive vertical measurements extend down the page. Thus, for margins closer to the left and top edges of the page than 1 inch, the corresponding parameters,

i.e., \evensidemargin, \oddsidemargin, \topmargin, can

be set to negative values.

To change the margins of a document within the document, modifying the parameters shown in Figure C.3 will not work. They can only be changed in the preamble of the document, i.e., before the \begin{document} statement. To adjust the margins within a document we define an environment:

\newenvironment{changemargin}[2]{%
 \begin{list}{}{%
  \setlength{\topsep}{0pt}%
  \setlength{\leftmargin}{#1}%
  \setlength{\rightmargin}{#2}%
  \setlength{\listparindent}{\parindent}%
  \setlength{\itemindent}{\parindent}%
  \setlength{\parsep}{0pt plus 1pt}%
 }%
\item[]}{\end{list}}

This environment takes two arguments, and will indent the left and right margins by their values, respectively. Negative values will cause the margins to be narrowed, so

\begin{changemargin}{-1cm}{-1cm} narrows the left and

right margins by 1cm. (Note that the value given for \parsep is the default for L^{A}TEX; what should be there is “the value in the surrounding text”.)

72 Finding the width of a letter, word, or phrase

Put the word in a box, and measure the width of the box. For example,

\newdimen\stringwidth
\setbox0=\hbox{hi}
\stringwidth=\wd0

Note that if the quantity in the \hbox is a phrase, the actual measurement only approximates the width that the phrase will occupy in running text, since the inter-word glue can be adjusted in paragraph mode.

The same sort of thing is expressed in L^{A}TEX by:

\newlength{\gnat}
\settowidth{\gnat}{\textbf{small}}

This sets the value of the length command \gnat to the width of “small” in emboldened text.

73 Excluding blocks of text from the .dvi file

Rainer Schöpf’s verbatim.sty provides a comment environment which excludes everything between \begin{comment} and \end{comment}. This package is available as part of

macros/latex/packages/tools

A more general environment for doing the job is Victor Eijkhout’s comment.sty, which lets you define environments for inclusion or exclusion in a document, thus offering a primitive configuration structure. It is available from the CTAN sites in

macros/latex209/contrib/misc/comment.sty

74 Defining a new log-like function in L^{A}TEX

Use the \mathop command, as in:

\newcommand{\diag}{\mathop{\rm diag}}

Subscripts and superscripts on \diag will be placed exactly as they are on \lim. If you want your subscripts and superscripts always placed to the right, do:

\newcommand{\diag}{\mathop{\rm diag}\nolimits}

This works in L^{A}TEX 2.09 and in L^{A}TEX 2ε, but not under NFSS alone (see question 81). There is an improved method for doing this in L^{A}TEX 2ε in the works; in the interim, one can use the the \operatorname command of amstex.sty (which is part of the AMS-L^{A}TEX package: fonts/ams/amslatex).

(It should be noted that “log-like” was reportedly a joke on Lamport’s part; it is of course clear what was meant.)

75 Typesetting all those TEX-related logos

Knuth was making a particular point about the capabilities of TEX when he defined the logo. Unfortunately, many believe, he thereby opened floodgates to give the world logos such as AMS-TEX, PICTEX, BibTEX, and so on. Lamport invented L^{A}TEX, and marketing input led to the current logo L^{A}TEX 2ε.

The common people don’t have to follow this stuff wherever it goes, but, for those who insist, a large collection of lo-

gos is defined in macros/eplain/texnames.sty; the  -

logo can be set in fonts that L^{A}TEX 2ε knows about (so that it scales with the surrounding text) using the package macros/

latex/contrib/supported/mflogo

For those who don’t wish to acquire the ‘proper’ logos, the canonical thing to do is to say AMS-\TeX{} (AMS-TEX) for AMS-TEX, Pic\TeX{} (PicTEX) for PICTEX, Bib\TeX{} (BibTEX) for BibTEX, and so on.

O Things are Going Wrong...

76 Weird hyphenation of words

You may have a version mismatch problem. The hyphenation system changed between version 2.9 and 3.0. If you are using (plain) TEX version 3.0 or later, make sure your plain.tex file has a version number which is at least 3.0; if you are using L^{A}TEX 2.09 (the last version of which was released on 25 March 1992), you need to upgrade to L^{A}TEX 2ε, since the sources of L^{A}TEX 2.09 are no longer publicly available.

If you’re using L^{A}TEX 2ε, the problem probably arises from your lthyphen.cfg file, which has to be created if you’re using a multi-lingual version.

For the curious, here’s what happened: before TEX 3.0 the hyphenation algorithm would not break a word if the part before the break was not at least two characters long, and the part after the break at least three characters long. Starting with version

3.0 the parameters \lefthyphenmin and \righthyphenmin

control the length of these fragments. These are set to 2 and 3, respectively, in the new plain and lplain formats. They can be set to any value, of course, but if \lefthyphenmin + \righthyphenmin is greater than 62, all hyphenation is suppressed.

77 (Merely) peculiar hyphenation

You may have found that TEX’s famed automatic word-division does not produce the break-points recommended by your dictionary. This may be because TEX is set up for American English, whose rules for word division (as specified, for example, in Webster’s Dictionary) are completely different from the British ones (as specified, for example, in the Oxford Dictionaries). This problem is being addressed by the UK TEX User community (see Baskerville, issue 4.4) but an entirely satisfactory solution will take time. An interim hyphenation file is available

in language/english/ukhyph.tex

78 Enlarging TEX

People sometimes get messages saying ‘memory capacity exceeded’. Most of the time this error can be fixed without enlarging TEX. The most common causes are unmatched braces, extra-long lines, and poorly-written macros. Extra-long lines are often introduced when files are transferred incorrectly between operating systems, and line-endings are not preserved properly (the tell-tale sign of an extra-long line error is when the complaint is that the ‘buf_size’ has overflowed).

If you really need to extend your TEX’s capacity, the proper method depends on your installation. In the purest form, you change the parameters in module 11 of the WEB source. In less pure forms, you might need to modify a change file, or perhaps change some environment variables; emTEX allows you to adjust the memory allocation criteria on the command line. Consult the documentation that came with your implementation.

79 Moving tables and figures in L^{A}TEX

Tables and figures have a tendency to surprise, by floating away from where they were specified to appear. This is in fact perfectly ordinary document design; any professional typesetting package will float figures and tables to where they’ll fit without violating the certain typographic rules. Even if you use the placement specifier h for ‘here’, the figure or table will not be printed ‘here’ if doing so would break the rules; the rules themselves are pretty simple, and are given on page 198, section C.9 of the L^{A}TEX manual. In the worst case, L^{A}TEX’s rules can cause the floating items to pile up to the extent that you get an error message saying “Too many unprocessed floats”; this means that the limited set of registers in which L^{A}TEX stores floating items is full. What follows is a simple checklist of things to do to solve these problems (the checklist talks throughout about figures, but applies equally well to tables).

• Are the placement parameters on your figures right? The default (tbp) is reasonable; you should never simply say ‘h’, for example, since that says “if it can’t go here, it can’t go anywhere”, and as a result all subsequent floats pile up behind it.

• Can you perhaps prevent your figures from floating by adjusting L^{A}TEX’s placement parameters? Again, the defaults are reasonable, but can be overridden in case of problems. The parameters are described on pages 199– 200, section C.9 of the L^{A}TEX manual.

• Are there places in your document where you could ‘naturally’ put a \clearpage command? If so, do: the backlog of floats is cleared after a \clearpage command. (Note that the \chapter command implicitly executes \clearpage, so you can’t float past the end of a chapter.)

• Have a look at the L^{A}TEX 2ε afterpage package (part

of macros/latex/packages/tools). Its documenta-

tion gives as an example the idea of putting \clearpage after the current page (where it will clear the backlog, but not cause an ugly gap in your text), but also admits that the package is somewhat fragile.

• As a last resort, try the package macros/latex209/

     contrib/misc/morefloats.sty; this ‘simply’ in-

creases the number of floating inserts that L^{A}TEX can handle at one time (from 18 to 36), but that may suit your needs.

• If you actually wanted all your figures to float to the end (e.g., for submitting a draft copy of a paper), don’t rely on L^{A}TEX to do it for you: get the endfloat pack-

     age to do the job for you (macros/latex/contrib/
     supported/endfloat).

80 \pagestyle{empty} on first page in L^{A}TEX

If you use \pagestyle{empty}, but the first page is numbered anyway, you are probably using the \maketitle command too. This is not a bug but a feature! The standard L^{A}TEX styles are written so that initial pages (pages containing a \maketitle, \part, or \chapter) have a different page style from the rest of the document. Hence, the commands internally issue \thispagestyle{plain}. This is usually not acceptable behaviour if the surrounding page style is ‘empty’.

Possible workarounds include:

• Put \thispagestyle{empty} immediately after the \maketitle command, with no blank line between them.

• Use fancyheadings.sty, which allows you to customise the style for initial pages independently of that for body pages. It is available in macros/latex/

     contrib/other/fancyheadings

81 Odd behaviour of \rm, \bf, etc.

If commands such as \rm and \bf have suddenly stopped working in L^{A}TEX in the way that you expect, it is likely that your system administrator has installed a version of L^{A}TEX 2.09 with NFSS (see question 94). Complain loudly; ask your system administrator to replace this version with L^{A}TEX 2ε (see question 95), in which commands such as \rm and \bf work just as before if you are using one of the standard classes—article, report and book (among others). In the meantime, use the option oldlfont.sty, which should have been installed at the same time as NFSS.

82 Old L^{A}TEX font references such as \tenrm

L^{A}TEX 2.09 defined a large set of commands for access to the fonts that it had built in to itself. For example, various flavours

of cmr could be found as \fivrm, \sixrm, \sevrm, \egtrm,
\ninrm, \tenrm, \elvrm, \twlrm, \frtnrm, \svtnrm,
\twtyrm and \twfvrm.

These commands were never documented, but certain packages nevertheless used them to achieve effects they needed.

Since the commands weren’t public, they weren’t included in L^{A}TEX 2ε; to use the unconverted L^{A}TEX 2.09 packages under L^{A}TEX 2ε, you need also to include the package rawfonts.sty (which is part of the L^{A}TEX 2ε distribution).

83 Missing symbols

If some symbols, such as \Box and \lhd, no longer appear to exist, then your system administrator has probably upgraded your version of L^{A}TEX to either NFSS (see question 94) or L^{A}TEX 2ε (see question 95). In the former case, use oldlfont.sty, as in the question 81. In the latter, use the package latexsym, which is part of the standard L^{A}TEX 2ε distribution, or the package amsfonts, if it is available.

84 L^{A}TEX gets cross-references wrong

Sometimes, however many times you run L^{A}TEX, the crossreferences are just wrong. Remember that the \label command must come after the \caption command, or be part of it. For example,

\begin{figure}          \begin{figure}
\caption{A Figure} or \caption{A Figure%
\label{fig}                   \label{fig}}
\end{figure}            \end{figure}

85 \@ and @ in macro names

A common source of problems in a L^{A}TEX document is the diagnostic about the appearance of the command \@, or about other commands containing the character @. The most common complaint is “You can’t use ‘\spacefactor’ in vertical mode”, but other similar ones occur.

Such problems are usually caused by including a L^{A}TEX 2ε class or package file into a L^{A}TEX document by some means

other than \documentclass or \usepackage. L^{A}TEX defines

internal commands whose names contain the character @; this enables it to avoid clashes between its internal names and names that we would normally use in our documents. In order that these commands may work at all, \documentclass and \usepackage play around with the meaning of @. Solve this problem by using the correct command to include the file.

But, you will say, “The L^{A}TEX Companion tells me to use commands containing @!”

Indeed; for example, there’s a lengthy section about \@startsection and how to use it to control the appearance of section titles. Page 15 of The Companion explains this; and suggests that you make such changes in the document pream-

ble, between \makeatletter and \makeatother. So the def-

inition of \subsection on page 26 could be:

  \makeatletter
  \renewcommand{\subsection}{\@startsection
    {subsection}%                    % name
    ...
    {\normalfont\normalsize\itshape}}% style
  \makeatother

86 Where are the msx and msy fonts?

The msx and msy fonts were designed by the American Mathematical Society in the very early days of TEX, for use in typesetting papers for mathematical journals. They were designed using the ‘old’ , which wasn’t portable and is no longer available; for a long time they were only available in 300dpi versions which only imperfectly matched modern printers. The AMS has now redesigned the fonts, using the current version of , and the new versions are called the msa and msb families; they are available from fonts/ams/

                                                     amsfonts/sources/symbols

Nevertheless, msx and msy continue to turn up to plague us. There are, of course, still sites that haven’t got around to upgrading; but, even if everyone upgraded, there would still be the problem of old documents that specify them.

If you have a .tex source that requests msx and msy, the best technique is to edit it so that it requests msa and msb (you only need to change the single letter in the font names).

If you have a dvi file that requests the fonts, there is a package of virtual fonts (see question 24) to map the old to the new

                                                     series; it’s available in fonts/vf-files/msx2msa

87 Where are the am fonts?

One still occasionally comes across a request for the am series of fonts. The initials stood for ‘Almost [Computer] Modern’, and they were the predecessors of the Computer Modern fonts that we all know and love (or hate)^{10}. There’s not a lot one can do with these fonts; they are (as their name implies) almost (but not quite) the same as the cm series; if you’re faced with a document that requests them, all you can reasonably do is to edit the document. The appearance of dvi files that request them is sufficiently rare that no-one has undertaken the mammoth task of creating a translation of them by means of virtual fonts; however, most drivers let you have a configuration file in which you can specify font substitutions. If you specify that every am font should be replaced by its corresponding cm font, the output should be almost correct.

88 ‘String too long’ in BibTEX

The BibTEX diagnostic “Warning–you’ve exceeded 1000, the global-string-size, for entry foo” is not one that you can hope to avoid by altering the BibTEX style in a simple way — BibTEX itself needs recompiling to increase its limit on string sizes (which is often not practical, and is never desirable). You must therefore address the problem by changing your bibliography database.

The problem usually arises from a very large abstract or annotation included in the database. The only way forward is to take the entry out of the database, so that you don’t encounter BibTEX’s limit, but you may need to retain the entry because it will be included in the typeset. In such cases, put the body of the entry in a separate file:

                                                       @article{long.boring,
                                                         author =    "Fred Verbose",
                                                         ...
                                                         abstract =  "{\input{abstracts/long.tex}}"
                                                       }

In this way, you arrange that all BibTEX has to deal with is the file name, though it will tell TEX (when appropriate) to include all the long text.

10The fonts acquired their label ‘Almost’ following the realisation that their first implementation in 79 still wasn’t quite right; Knuth’s original intention had been that they were the final answer

P Why does it do that?

89 Why does it ignore paragraph parameters?

When TEX is laying out text, it doesn’t work from word to word, or from line to line; the smallest complete unit it formats is the paragraph. The paragraph is laid down in a buffer, as it appears, and isn’t touched further until the end-paragraph marker is processed. It’s at this point that the paragraph parameters have effect; and it’s because of this sequence that one often makes mistakes that lead to the paragraph parameters not doing what one would have hoped (or expected).

Consider the following sequence of L^{A}TEX:

{\raggedright % declaration for ragged text
Here's text to be ranged left in our output,
but it's the only such paragraph, so we now
end the group.}
Here's more that needn't be ragged...

TEX will open a group, and set the ragged-setting parameters within that group; it will then save a couple of sentences of text and close the group (thus restoring the previous value of the ragged-setting parameters). Then it encounters a blank line, which it knows to treat as a \par token, so it typesets the two sentences; but because the enclosing group has now been closed, the parameter settings have been lost, and the paragraph will be typeset normally.

The solution is simple: close the paragraph inside the group, so that the setting parameters remain in place. An appropriate way of doing that is to replace the last three lines above with:

end the group.\par}
Here's more that needn't be ragged...

In this way, the paragraph is completed while the setting parameters are still in force within the enclosing group.

90 What’s the reason for ‘protection’?

Sometimes L^{A}TEX saves data it will reread later. These data are often the argument of some command; they are the socalled moving arguments. (‘Moving’ because data are moved around.) Places to look for are all arguments that may go into table of contents, list of figures, etc.; namely, data that are written to an auxiliary file and read in later. Other places are those data that might appear in head- or footlines. Section headers and figure captions are the most prominent examples; there’s a complete list in Lamport’s book (see question 17).

What’s going on really, behind the scenes? The commands in the moving arguments are already expanded to their internal structure during the process of saving. Sometimes this expansion results in invalid TEX code when processed again. “\protect\cmd” tells L^{A}TEX to save \cmd as \cmd, without expansion.

What is a ‘fragile command’? It’s a command that expands into illegal TEX code during the save process.

What is a ‘robust command’? It’s a command that expands into legal TEX code during the save process.

No-one (of course) likes this situation; the L^{A}TEX3 team have removed the need for protection of some things in the production of L^{A}TEX 2ε, but the techniques available to them within current L^{A}TEX mean that this is an expensive exercise. It remains a long-term aim of the team to remove all need for these things.

91 Why doesn’t \verb work within...?

The L^{A}TEX verbatim commands work by changing category codes. Knuth says of this sort of thing “Some care is needed to get the timing right...”, since once the category code has been assigned to a character, it doesn’t change. So \verb has to assume that it is getting the first look at its parameter text; if it isn’t, TEX has already assigned category codes so that \verb doesn’t have a chance. For example:

    \verb+\error+

will work (typesetting ‘\error’), but

    \newcommand{\unbrace}[1]{#1}
    \unbrace{\verb+\error+}

will not (it will attempt to execute \error).

This is why the L^{A}TEX book insists that verbatim commands must not appear in the argument of any other command; they aren’t just fragile, they’re quite unusable in any command parameter, regardless of \protection (see question 90).

92 Case-changing oddities

TEX provides two primitive commands \uppercase and \lowercase to change the case of text; they’re not much used, but are capable creating confusion.

The two commands do not expand the text that is their parameter — the result of \uppercase{abc} is ‘ABC’, but \uppercase{\abc} is always ‘\abc’, whatever the meaning of \abc. The commands are simply interpreting a table of equivalences between upper- and lowercase characters. They have (for example) no mathematical sense, and

  \uppercase{About $y=f(x)$}

will produce

  ABOUT $Y=F(X)$

which is probably not what is wanted.

The same sort of problem can arise with L^{A}TEX environment names, on occasions when L^{A}TEX employs \uppercase implicitly. For example, the standard classes use \uppercase to set the marks according to the arguments of \chapter and \section commands, so that a chapter whose title contains a small table expressed in terms of

\begin{tabular} ... \end{tabular}, will set a mark con-

taining \begin{TABULAR}, etc., which is an environment that doesn’t exist.

93 Why are # signs doubled in macros?

The way to think of this is that ## gets replaced by # in just the same way that #1 gets replaced by ‘whatever is the first argument’.

So if you define a macro and use it as:

  \def\a#1{...#1...#1...#1...}  \a{b}

the macro expansion produces ‘...b...b...b...’, which people find normal. However, if we now fill in the ‘...’:

\def\a#1{----#1---\def\x #1{xxx#1}}

\a{b} will expand to ‘---b---\def\x b{xxxb}’. This defines \x to be a macro delimited by b, and taking no arguments, which people may find strange, even though it is just a specialisation of the example above. If you want \a to define \x to be a macro with one argument, you need to write:

\def\a#1{---#1---\def\x ##1{xxx##1}}

and \a{b} will expand to ‘---b---\def\x #1{xxx#1}’, because #1 gets replaced by ‘b’ and ## gets replaced by #.

To nest a definition inside a definition inside a definition then you need ####1, as at each stage ## is replaced by #. At the next level you need 8 #s each time, and so on.

Q Recent Developments

94 The New Font Selection Scheme (NFSS)

NFSS was an extension to L^{A}TEX written by Frank Mittelbach and Rainer Schöpf. It is described in TUGboat, 1989 10(2). In traditional typesetting, fonts are described by four parameters: the family (e.g., computer modern), the series (i.e., the weight and width of the font, such as light or bold), the shape (e.g., italic), and the size. NFSS is a mechanism allowing the user to change any of these independently. NFSS makes it relatively easy to use nonstandard fonts such as the PostScript ones with L^{A}TEX, and easy to change maths fonts. It also allows dynamic loading of fonts at runtime (i.e., not when the format file is created).

With the demise of L^{A}TEX 2.09 as supported software, the label ‘NFSS’ has become somewhat misleading, as there’s no ‘old’ scheme with which to contrast it — L^{A}TEX has incorporated the NFSS.

95 L^{A}TEX 2ε (the new standard L^{A}TEX)

L^{A}TEX 2ε is a new version of the L^{A}TEX package, prepared and supported by the L^{A}TEX3 project team. It moved out of its test phase in June 1994, and is now the standard L^{A}TEX; L^{A}TEX 2.09 is no longer supported.

L^{A}TEX 2ε is upwardly compatible with L^{A}TEX 2.09, but has new features, including:

• NFSS (see question 94) is now standard.

• SLiTEX is consolidated into it, so that there is no longer a need for a separate format.

• Better control of floating environments, such as figures.

• There is a documented interface for package and class writers (though not yet for designers).

• Enhanced box commands, e.g., options to specify the height of a minipage.

• \ref can be used in \caption without being protected.

• \newcommand can define commands with one optional argument.

• A standard package for colour and graphics inclusion.

Since L^{A}TEX 2ε is supported, you can report bugs or problems with it by typing ‘latex latexbug’ and sending the re-

port thus generated to latex-bugs@uni-mainz.de.

96 The L^{A}TEX3 project

The L^{A}TEX3 project team is a small group of volunteers whose aim is to produce a major new document processing system based on the principles pioneered by Leslie Lamport in the current L^{A}TEX. It will remain freely available and it will be fully documented at all levels.

The L^{A}TEX3 team has already delivered its first product, L^{A}TEX 2ε (see question 95), a macro package based on Lamport’s original code, but modified to be more readily supportable than was Lamport’s.

97 Using TEX to typeset from SGML files

Jonathan Fine (J.Fine@pmms.cam.ac.uk) is developing software which will allow TEX to typeset directly from SGML files. The work is producing a macro package that directly interprets an SGML source file, as opposed to the conversion programs outlined earlier (see question 45).

The work is beginning to show fruits of success, and an announcement will be made early in 1995.

98 The Omega project

Omega (Ω) is a program built on top of TEX which works internally with 16-bit characters (Unicode); this allows it to work with most scripts in the world without any complications of coding schemes. Omega also has a powerful concept of input and output filters to allow the user to work with existing transliteration schemes, etc. Omega is an ongoing project by John Plaice (plaice@ift.ulaval.ca) and Yannis Har-

alambous (Yannis.Haralambous@univ-lille1.fr). Con-

tact Yannis to get on the discussion list.

99 The N_{T}S project

The NTS project first saw the light of day at the Hamburg meeting of DANTE during 1992, as a response to an aspiration to produce something even better than TEX. The project is not simply enhancing TEX, for two reasons: first, that TEX itself has been frozen by Knuth (see question 13), and second, even if they were allowed to develop the program, some members of the NTS team feel that TEX in its present form is simply unsuited to further development. While all those involved in the project are involved with, and committed to, TEX, they recognise that the end product may very well have little in common with TEX other than its philosophy.

Initially, and despite the reservations expressed at the inaugural meeting, the group is concentrating on extending TEX per se: members are implementing extensions and enhancements to TEX through the standard medium of a change-file. These extensions and enhancements, together with TEX proper, will form a system called ε-TEX, which will be 100% compatible with TEX; furthermore, it will be possible during format creation to construct a format that is TEX: no extensions or enhancements will be present.

The final aim of the project will be to produce an entirely new typesetting system, building on the experience gained in the earlier phases. This system is intended to provide a stable basis for typesetting in the future, in the way that TEX has since it was first offered to the world.

R Perhaps There isn’t an Answer

100 What to do if you find a bug

For a start, make entirely sure you have found a bug. Doublecheck with books about TEX, L^{A}TEX, or whatever you’re using; compare what you’re seeing against the other answers above; ask every possible person you know who has any TEX-related expertise. The reasons for all this caution are various.

If you’ve found a bug in TEX itself, you’re a rare animal indeed. Don Knuth is so sure of the quality of his code that he offers real money prizes to finders of bugs; the cheques he writes are such rare items that they are seldom cashed. If you think you have found a genuine fault in TEX itself (or , or the CM fonts, or the TEXbook), don’t immediately write to Knuth, however. He only looks at bugs once or twice a year, and even then only after they are agreed as bugs by a small vetting team. In the first instance, contact Barbara Beeton at the AMS (bnb@math.ams.org), or contact TUG.

If you’ve found a bug in L^{A}TEX 2ε, you should submit its details to the L^{A}TEX3 team. To do this, you should use L^{A}TEX to process the file latexbug.tex, which is part of the L^{A}TEX 2ε distribution. The process will give you instructions about what to do with your bug report. Please be sparing of the team’s time; they’re doing work for the good of the whole L^{A}TEX community, and any time they spend tracking down non-bugs is time not available to write or debug new code.

If you’ve found a bug in L^{A}TEX2.09, or some other such unsupported software, there’s not a lot you can do about it. You may find help or de facto support on a newsgroup such as comp. tex.tex or on a mailing list such as uktex@tex.ac.uk, but posting non-bugs to any of these forums can lay you open to ridicule! Otherwise you need to go out and find yourself a willing TEX-consultant^{11}.

11TUG maintains a register of TEX consultants; UKTUG is developing one